### v2

# Network Formation and Dynamics Among Multi-LLMs

Abstract.
Social networks shape opinions, behaviors, and information dissemination in human societies. As large language models (LLMs) increasingly integrate into social and professional environments, understanding their behavior within the context of social interactions and networks becomes essential. Our study analyzes LLMs’ network formation behavior to examine whether the dynamics of multiple LLMs are similar to or different from human social dynamics. We observe that LLMs exhibit key social network principles, including preferential attachment, triadic closure, homophily, community structure, and the small-world phenomenon, when asked about their preferences in network formation. We also investigate LLMs’ decision-making based on real-world networks, revealing that triadic closure and homophily have a stronger influence than preferential attachment and that LLMs perform well in network formation predictions. Overall, our study opens up new possibilities for using LLMs in network science research and helps develop socially aware LLMs by shedding light on their network formation behaviors and exploring their impacts on social dynamics.
large language models, networks, network formation, link prediction, multi-agent systems
conference: ; ;
1. Introduction
Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023) and LLaMA2 (Touvron et al., 2023), have shown promising developments in AI techniques and their integration into real-life applications. It is thus crucial to comprehend AI actions to ensure they align with human expectations, mitigate potential risks, and maximize their benefits. Misaligned AI actions may lead to unintended consequences, such as biased decision-making, fairness issues, and the miscoordinative or non-cooperative behavior (Rahwan et al., 2019). Recently, researchers have started to apply methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods, to study LLMs. These methods help uncover the generative capabilities of LLMs and their interpretability (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
In human societies, social networks play a crucial role in shaping individual behaviors, preferences, and connections, as well as influencing the diffusion of information and norms across communities (Rogers, 2003; Bakshy et al., 2012; Fowler and Christakis, 2008; Banerjee et al., 2013; Yuan et al., 2018). LLMs have shown great potential in social contexts, notably as intelligent personal assistants facilitating social and business interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)). However, less is known about how LLMs’ innate behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023). This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the intricate networks of human relationships, which is a fundamental aspect of social systems.
Our study thus explores LLMs’ behaviors and preferences in network formation in both synthetic and real-world social networks. By analyzing interactions between multiple LLMs (or multi-LLMs), we aim to understand the implications of LLMs representing humans in social and business settings. Specifically, we examine micro-level social network properties including preferential attachment (Barabási and Albert, 1999), triadic closure (Granovetter, 1973), and homophily (McPherson et al., 2001), as well as macro-level properties including community structure (Newman and Girvan, 2004), and the small-world phenomenon (Kleinberg, 2000; Watts and Strogatz, 1998).
In synthetic networks, our findings reveal that LLMs can generate networks exhibiting preferential attachment and following scale-free degree distributions. Moreover, LLMs demonstrate strong tendencies toward triadic closure and exhibit tendency of homophily. These tendencies further explain the emergence of community structures with high modularity in networks formed by multiple LLMs. Additionally, we show that LLM agents exhibit dynamics consistent with small-world phenomena akin to those observed in the Watts-Strogatz model (Watts and Strogatz, 1998).
Next, we conduct a comparative analysis of these principles to identify the key factors driving network dynamics in the context of real-world network formation scenarios. To do so, we consider real-world networks and ask each LLM to simulate a node in the network. We provide information on their network structure and ask for their decisions for the next link formation or dissolution step. We approach this problem as a discrete choice modeling problem (Liben-Nowell and Kleinberg, 2003; Overgoor et al., 2019). In this real-world setting, we find that LLMs’ choices are primarily weighted on homophily and triadic closure, compared to preferential attachment, and that LLMs substantially exceed random guessing when predicting new links. Additionally, the resulting choices of LLM agents strengthen the network’s community structure.
Our study contributes to the literature on LLMs in several ways. First, it advances the discussion on aligning LLMs with human social preferences, which has potential applications in developing socially aware LLMs for real-life interactions. Insights into LLMs’ innate behaviors in network formation can inform the design of LLMs for effective collaboration in social and professional settings. Second, our work has direct applications to synthetic dataset generation, a problem important for the graph learning community. By controlling for a few parameters, such as the temperature, our work can be utilized to generate a variety of datasets, as well as generate synthetic datasets from real-world data to protect privacy. Finally, our study could contribute to understanding the impact of LLMs on social dynamics and norms, such as their role in shaping online communities’ interaction patterns or influencing the adoption of new communication norms in professional settings.
    # Task     Your task is to select a set of people to be friends with.     # Profile     Your profile is given below after chevrons:     <PROFILE>$F(\{i_{t}\})$</PROFILE>     # Candidate Profiles     The candidate profiles to be friends with are given below after chevrons:     <PROFILES>$F(A_{t})$</PROFILES>     # Output     The output should be given a list of JSON objects with the following structure     [         {{             "name" : name of the person you selected,             "reason" : reason for selecting the person         }}, ...     ]     # Notes     - The output must be a list of JSON objects ranked in the order of preference.     - You can make at most $\delta$ selections.
ALGORITHM 1 Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$ formatted as a Python f-string.
(a) Structure-based
(b) Degree-based
Figure 1. Networks with $n=200$ nodes created based on Principle 1 for various temperatures. For the structure-based networks (Figure 1(a)), we provide the multi-LLMs with the neighborhoods $\{N_{j,t}:j\in i_{t}\}$, and for the degree-based networks (Figure 1(b)), we provide the degrees $\{d_{j,t}:j\in i_{t}\}$ of the nodes $j\in i_{t}$. We also plot the degree distribution and compare it against a BA network with $n=200$ nodes. Top: In Figure 1(a), we observe that the resulting networks follow a power-law degree distribution ($P>0.5$, Bootstrap hypothesis test). When the temperature is 1.5, the resulting network resembles the BA model ($P<0.001$, K-S test). Bottom: When the degree information is provided only (Figure 1(b)), the resulting networks correspond to unrealistic star-like formations that are far from a scale-free network and resemble more a core-periphery network.
2. Results
Our study employs the state-of-the-art LLMs, such as GPT-3.5 and GPT-4 (OpenAI, 2023) with the OpenAI API [^1] For GPT-3.5 we use the gpt-3.5-turbo model provided by OpenAI, and for GPT-4 we use the gpt-4-1106-preview model.. We constructed multiple LLMs using multiple independent conservation threads where each thread represents one multi-LLM. We prompt each LLM by providing information about other LLMs’ network information as well as decisions they are faced with. We test both state-of-the-art GPT-based models: for the results of Sections 2.2 and 2.1, we use GPT-3.5 agents, and in Section 2.3, we use GPT-4 agents. We obtained similar results using GPT-3.5 and GPT-4. Overall, we adopted a standard framework for multi-LLMs but underscored the multi-agent nature of the design in our setting (Leng et al., 2023; Akata et al., 2023; Brookins and DeBacker, 2023; Aher et al., 2023; De Marzo et al., 2023). We evaluate the robustness of our results by testing various temperature settings in LLMs, which control the degree of variation of the model output. All experiments employ a dynamic network setup. We also employ an interview-like approach to inquire about the rationale behind the decisions of LLMs. The details of our methods are presented in Methods.
In this section, we first discuss three major micro-level network principles: preferential attachment, triadic closure, and homophily. We then examine two macro-level network principles: community structure and small-world phenomenon. Subsequently, we test the strength of these principles when LLM agents face real-world networks and compare the preferences in network formation tasks between LLMs and humans.
2.1. Micro-Level Properties
Principle 1: Preferential Attachment
Preferential attachment is a well-known mechanism in network science that explains how certain nodes in a network become more connected over time, leading to the emergence of highly connected vertices and scale-free distribution of degree (Barabási and Albert, 1999; Bianconi and Barabási, 2001).
We examine whether LLMs exhibit the preferential attachment mechanism in the following way: First, we start with an empty network and add nodes at each time step $t$. For each new time step, we add a new node $i_{t}$, providing information on the network neighbors of all existing nodes. [^2] See Appendix E, for information about how the features are represented in the prompt. We chose to generate networks with $n=200$ nodes, as this size is sufficiently large to allow for the observation of degree distribution patterns.
We plot one of the examples of resulting networks under each temperature in Figure 1(a). As observed in all temperature setups, we observe a few nodes that have many connections, and the majority of the notes have one or few connections, which renders the degree distribution scale-free, i.e., the probability $\pi(d)$ a node has degree $d$ is given by
$\pi(d)\propto d^{-\gamma},\quad\text{where }\gamma>1.$
Furthermore, we observe mode “super nodes” with many connections as the temperature increases.
We also compare with the Barabási–Albert (BA) model (Barabási and Albert, 1999) where each incoming node $t$, the probability $p_{i,t}$ of it connecting to node $i$ is proportional to ${d_{i,t}}$, where $d_{i,t}$ is the degree of node $i$ at time step $t$, namely,
$p_{i,t}=\frac{d_{i,t}}{\sum_{j=1}^{t}d_{j,t}}.$
For temperatures of 0.5 and 1.0, we find that the estimated exponent values are $\hat{\gamma}=1.50\pm 0.23$ and $1.92\pm 0.24$, respectively. However, the degree distributions do not resemble the degree distribution of the generated BA network with an expected exponent $\gamma=3$. This is supported by the $P$-values from a two-sample Kolmogorov-Smirnov (K-S) test, which are $P<0.001$ and $P<0.05$, respectively when compared to the empirical degree distribution of the BA network [^3] As shown by (Barabási, 2013), the BA model follows a scale-free distribution of the form $\pi_{\mathrm{BA}}(d)\propto d^{-3}$ for sufficiently large $n$. However, because our $n$ is small, so that the prompt does not exceed the token limit for the GPT-3.5 model, we directly compared it to a sample from the BA model (the generated network has $\hat{\gamma}_{\mathrm{BA}}=2.24\pm 0.24$), instead of comparing with the limiting distribution $\pi_{\mathrm{BA}}(d)$..
Moreover, for all temperatures, the resulting degree distribution is scale-free ($P>0.5$, bootstrapping hypothesis test comparing with the fitted scale-free distribution).
We also examine the network formation when only degree information (rather than network structure) is provided, in a similar spirit of (De Marzo et al., 2023). We present the results in Figure 1(b). We observe “star-like” formations around one node. The degree distributions differ from scale-free distributions produced by statistical models such as the BA model ($P<0.001$, two-sample K-S test compared to the empirical distribution of the generated BA network). They are closer to a core-periphery network (Borgatti and Everett, 2000).
Finally, we employ text analysis of reasoning provided by the agents for the results of Figure 1(a) [^4] We do not provide the distribution of reasons for the results of Figure 1(b) as all of the LLM’s decisions are made given the degree information.. We observe two main categories of reasons (see Appendix C) for the fraction under each temperature):
  • Well-Connectedness: it corresponds to reasons related to the person being well-connected and having many friends.
  • Mutual Friends: it is related to having many friends in common with the person so that the person can introduce them to their friends.
The well-connectedness reason becomes less prevalent as the temperature increases, and the mutual friends’ reason becomes more prevalent. Note that when the temperature is 1.5, the two reasons have almost the same frequency.
Taken together, networks created by LLM systems, when given a network structure, demonstrate preferential attachment and exhibit scale-free degree distributions.
Principle 2: Triadic Closure
(a) Structure-Based
(b) Common-Neighbor-Based
Figure 2. Networks created based on Principle 2. Top: Figure 2(a) shows the resulting networks created by Principle 2 when the intersection of the neighborhoods of the query node and each alternative is provided. The node colors correspond to the stochastic block model (SBM) groups each node belongs to. The red edges correspond to the newly created edges between nodes of the red cluster, the blue edges correspond to the newly created edges between nodes of the blue cluster, and the orange edges correspond to edges created between the two clusters. The newly formed edges are drawn with a larger line width. We observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). Bottom: Figure 2(b) shows the same networks with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative. Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test).
The second micro-level principle we examine is triadic closure, which describes the phenomenon that individuals are likely to form connections with their friends’ friends, leading to closed triads within the network. This tendency enhances the network’s structure and cohesion, rooted in the likelihood of two nodes forming a direct connection if they share a common neighbor (Granovetter, 1973; Mosleh et al., 2024).
To test for triadic closure, we use an assortative stochastic block model (SBM) (Newman, 2003) to generate an initial seed network $G_{1}$ with $n$ nodes divided into two equal-sized clusters $A$ and $B$. Connections within each cluster have a probability of $0.5$, while connections across clusters have a probability of $0.1$. This setup reflects our intuition that nodes within the same cluster are more likely to form connections due to their higher number of common neighbors. We then iterate over each node $i$, providing features based on the intersection of the neighborhoods of all of $i$’s non-neighbors or the number of common friends between $i$ and any non-neighbor $j$. This modeling choice contrasts the results mentioned about preferential attachment, where the neighborhood information is provided and made to make the two settings consistent [^5] We want to note that performing the same experiment when we provide the neighbors instead of the common neighbors produces similar results..
We compare this with the random null model, which selects a node at random to connect to. We run ten simulations for $n=50$ nodes. [^6] For better visualization, we chose $n=50$ nodes instead of a large number of nodes such as $n=200$. Moreover, the results are statistically significant even with $n=50$ nodes.
To measure whether the principle of triadic closure exists, we employ the fraction $\hat{p}$ of newly added edges between nodes of the same community, i.e., the fraction of edges in $G_{T}\setminus G_{1}$ that have their ends both at either $A$ or $B$, namely
$\hat{p}=\frac{\left|\left\{\{i,j\}\in E(G_{T})\setminus E(G_{1}):y_{i}=y_{j}\right\}\right|}{\left|E(G_{T})\setminus E(G_{1})\right|},$
where $y_{i},y_{j}\in\{A,B\}$ are the community memberships of nodes $i$ and $j$ respectively, and $E(G_{T})\setminus E(G_{1})$ corresponds to the newly added edges.
Generally, a value of $\hat{p}>0.5$ indicates the tendency of triadic closure since an LLM agent, which is supposed to exhibit triadic closure given the common neighbors, would ideally create a link within the same cluster.
Additionally, we introduce the concept of marginal transitivity for the network, represented by $D$. This metric is defined as the change in the proportion of closed triangles to open triads in the network, from $G_{T}$ (the initial network) to $G_{1}$ (which corresponds to the SBM):
$D=3\times\frac{\text{\# triangles}(G_{T})}{\text{\# triads}(G_{T})}-3\times\frac{\text{\# triangles}(G_{1})}{\text{\# triads}(G_{1})}.$
We depict the sample outcomes in Figure 2(a), where the upper panel presents the results when the entire network structure is provided, whereas the lower panel only provides the number of common neighbors. Node colors correspond to their respective groups within the SBM. Red edges denote connections within the red cluster, blue edges within the blue cluster, and orange edges between the two clusters. Newly established edges are emphasized with a thicker line width.
We present the probabilities of network formation within a community in the last panels in each row. We find that the probability of forming a within-community tie across all temperatures is substantially larger than the random network formation scenario ($P<0.001$, t-test comparing to Random), suggesting LLMs’ tendency of triadic closure (since an LLM which is given the common neighbors as features should ideally create links within the same community).
Additionally, we observe a statistically significant increase in the marginal transitivity $D$ for all temperatures ($P<0.001$, t-test). Moreover, in all cases, the marginal transitivity $D$ is statistically significant compared to Random ($P<0.001$, t-test).
These results suggest that LLM agents exhibit the tendency of triadic closure. Similar to the approach in earlier reasoning analysis, we present the most common reasons provided by LLMs in Appendix C [^7] We do not provide the distribution of reasons for the results of Figure 2(b) as all of the LLM’s decisions are made given the number of common neighbors information.. We find that most of the agents base their decisions on the number of mutual friends.
Principle 3: Homophily
(a) Assortativity Coefficients
(b) Network Instances
(c) Community Structure
Figure 3. Top: Figure 3(a) shows the assortativity coefficients for each network attribute created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. We observe that for all temperatures, we have positive assortativity ($P<0.001$, t-test with 0), which indicates homophily. The assortativities are significantly different for all temperatures from random or heterophilous ($P<0.001$, t-test comparing with Random) Middle: Networks created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. The edge widths correspond to the number of common attributes shared between the nodes. We report the probability of forming a link as a function of the number of common attributes between two nodes. The node colors correspond to the communities found by running the Louvain algorithm. Bottom: Louvain modularity, number of communities, and community sizes of networks created based on Principle 3. Lower temperatures yield higher modularity ($P<0.001$, t-test) and fewer and larger communities ($P<0.001$, t-test). For smaller temperatures, the communities are more “well-separated”.
Homophily reflects the tendency for nodes with similar characteristics or attributes to form connections and associate with each other. This phenomenon is based on the principle that individuals in a network are more likely to connect with others who share similar traits, interests, or demographics (McPherson et al., 2001).
To test whether LLM agents exhibit homophily, we perform the following experiment: We generate nodes with randomly generated attributes regarding a hobby (randomly chosen among three hobbies), a favorite color (randomly chosen among three colors), and a location within the US (randomly chosen among three US locations) and provide the attributes of the other nodes and the node’s own attributes, and each node is tasked to form up to $\delta=5$ links with others. For each node $i$, we provide it with the features $x_{j}$ of all non-neighbors $j$ of $i$. The seed network is taken to be the empty graph.
We run ten simulations for networks with $n=50$ nodes and $\delta=5$. We compare with the Random null model, which chooses $\delta=5$ links at random.
To measure the presence of homophily, we measure the attribute assortativity coefficient of each of the three features (hobby, color, location). For a property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$, its assortativity coefficient $R$ is defined as
$R=\frac{\sum_{i=1}^{n}M_{ii}-\sum_{i=1}^{n}a_{i}b_{i}}{1-\sum_{i=1}^{n}a_{i}b_{i}},$
where $M$ is the mixing matrix with entries $M_{ij}$ corresponding to the fraction of edges where one end has value $P_{i}$, and the other end has value $P_{j}$, $a_{i}=\sum_{j=1}^{n}M_{ij}$ and $b_{i}=\sum_{j=1}^{n}M_{ji}$. The assortativity has values between $-1$ and $+1$, where a positive value denotes the tendency of the nodes to connect to other similar nodes (hence yielding a homophilous network), and a negative value corresponds to connecting to the tendency to connect to dissimilar nodes (which indicates heterophily).
In Figure 3, we observe a consistent pattern of positive assortativity across all temperature ranges, which is higher for lower temperatures. Moreover, the observed assortativity at all temperatures significantly deviates from Random, which has assortativity close to 0 ($P<0.001$, t-test comparing with Random). Taken together, we conclude homophily emerges for the networks generated by LLMs.
Moreover, we present the probability of each reason being mentioned in Appendix C. We find that shared hobbies appear to be much more important than shared location or favorite color. This is consistent with our intuition that, similarly to humans, LLMs care more about hobbies than other superficial attributes when forming social connections.
2.2. Macro-Level Principles
Principle 4: Community Structure
The community structure of networks refers to the organization of nodes or individuals within a network into distinct and densely interconnected groups or clusters (Newman and Girvan, 2004; Newman, 2006; Blondel et al., 2008; Clauset et al., 2004). Identifying community structures is crucial for understanding the overall dynamics of a network, as it reveals patterns of relationships and interactions that might not be apparent at the global level.
Both triadic closure and homophily contribute to the formation of community structures.By examining how these two factors contribute to network formation, we aim to gain insights into the underlying mechanisms driving community dynamics in LLM-generated networks. We employ the simulation results presented in Section 2.1 to determine whether community structure in networks generated by LLMs emerges from triadic closure or homophily.
First, we consider the networks generated in Figure 2. We examine how LLM agents’ choices strengthen the network’s community structure. Specifically, we leverage the fact that the SBM graph has a preexisting community structure and measure how the newly formed links reinforce such a structure. Visual inspection shows that the newly added links, represented by the bold edges, happen mostly within each cluster, reinforcing the community structure. This is further quantitatively verified by the fraction $\hat{p}$ newly created inter-community edges. We find that $\hat{p}$ is significantly higher than 0.5 ($P<0.001$, t-test comparing with 0.5). This indicates that most edges are within the same community, strengthening the community structure.
Next, we use modularity maximization (Blondel et al., 2008) to examine the community structure that arises due to homophily. Intuitively, the modularity measure compares the number of edges inside a community with the expected number of edges that one would find in the cluster if the network were a random network with the same number of nodes and where each node keeps its weighted degree, but edges are otherwise randomly attached according to the Chung-Lu model (Chung and Lu, 2004). This random null model implicitly assumes that each node can attach to another network node. Formally, for a graph with weights $w_{ij}$ on its edges, the weighted modularity $Q$ (Clauset et al., 2004) of a partition with $C$ communities given by
$Q=\sum_{c=1}^{C}\left[\frac{L_{c}}{W}-r\left(\frac{k_{c}}{2W}\right)^{2}\right],$
Here $W$ is the sum of all weights, $L_{c}$ is the total weight of intra-community links for community $c$, $k_{c}$ is the total weighted degree of the nodes in the community $c$ and $r$ is the resolution parameter which (we take $r=1$ from now on). A large positive value (such as greater than 0.5) of the modularity corresponds to evidence of community structure, corresponding to deviation from the random null model. Maximizing $Q$ is NP-Hard, we use the Louvain algorithm (Blondel et al., 2008) to get a lower bound on the maximum modularity.
In Figure 3(b), we illustrate the communities identified by the Louvain algorithm at various temperatures. Alongside, we present the average Louvain modularity values over ten simulations, with 95% confidence intervals (refer to Figure 3(c)). For the network’s weights, we use the number of common attributes shared between each pair of nodes. This is defined as $w_{ij}=\left|\left\{k:x_{i}^{(k)}=x_{j}^{(k)}\right\}\right|$ for each link $(i,j)$ in the final network. Here, $x_{i}^{(k)}$ and $x_{j}^{(k)}$ correspond to the $k$-th features of $x_{i}$ and $x_{j}$, respectively.
We observe that the Louvain partitions demonstrate community structure with positive modularity across all temperatures. This is confirmed by a t-test comparing the modularity to $Q=0$ of a random graph ($P<0.001$). The modularity is most pronounced at lower temperatures. As the temperature increases, the Louvain modularity also increases ($P<0.001$, t-test). Furthermore, lower temperatures result in fewer (and larger) communities on average, as identified by the Louvain algorithm ($P<0.001$, t-test). Additionally, communities tend to be visually more distinct at lower temperatures. We propose that higher temperatures reflect increased randomness in the decisions.
Our results demonstrate that community structures manifest in networks generated by LLMs. Specifically, these networks adhere to the principle of triadic closure, which further reinforces the formation of community structures. Additionally, LLM-generated networks exhibit homophily, facilitating the emergence of distinct communities. In Section 2.3, we further illustrate that community structures are reinforced in real-world networks, where LLMs consider a combination of factors, including homophily and triadic closure, in their decisions to form connections.
Principle 5: Small-World
(a) Network Instances for $\beta=0.25$
(b) Network Instances for $\beta=0.5$
(c) Network Instances for $\beta=0.75$
(d) Estimated $\hat{\beta}$ for $\beta=0.25$
(e) Estimated $\hat{\beta}$ for $\beta=0.5$
(f) Estimated $\hat{\beta}$ for $\beta=0.75$
Figure 4. Figures 4(a), 4(b) and 4(c): and distribution of reasoning created based on Principle 5 using the altered Watts-Strogatz Model for $n=50,k=5,\beta\in\{0.25,0.5,0.75\}$, together with plots of the average clustering coefficient $C$ and the average shortest path length $L$. The comparison is made with respect to a Watts-Strogatz graph with $n=50,k=5,\beta\in\{0.25,0.5,0.75\}$. In most cases, we observe strong evidence that $L\sim\log(n)$ and $C\sim 1/\log(n)$. Figures 4(d), 4(e) and 4(f): Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with rewiring probability $\hat{\beta}$. We report the $P$-values of the t-test comparing the average shortest path length of the LLM-generated networks and the average shortest path length of the Watts-Strogatz graphs with rewiring probability $\hat{\beta}$.
The small-world phenomenon is characterized by networks where nodes are interconnected in tight clusters, yet the average distance between any two nodes remains relatively short, typically scaling logarithmically with the network size (Watts and Strogatz, 1998; Kleinberg, 2000). This balance between high clustering and short path lengths characterizes small-world networks.
Formally, a small-world network is a network where the average shortest path length $L$ grows proportionally to the network size $n$, [^8] This follows the definition in (Jacobs and Watts, 2021). namely
$L\sim\log(n).$
Our analysis utilizes the Watts-Strogatz model (Watts and Strogatz, 1998) as a benchmark to investigate whether LLMs can generate networks exhibiting small-world characteristics. This model has a delicate balance between local clustering and short average path lengths: Nodes tend to form clusters or groups (triadic closure), exhibiting a high level of interconnectedness within these local neighborhoods, whereas at the same time, the existence of a few long-range connections ensures that the entire network is reachable with relatively few steps (Park et al., 2018; Lyu et al., 2022; Jahani et al., 2023).
We employ a modified version of the model, where edge rewiring is informed by LLM queries, based on the current network structure. The generation process is parametrized by the number of nodes ($n$), the number of long-range connections ($k$), and the rewiring probability ($\beta$). The algorithm proceeds as follows:
  1. (1)
    Similarly to Watts-Strogatz, we first create a ring network with $n$ nodes. After that, for each node $[n]$, we create $k$ edges where $k/2$ edges connect to its rightmost neighbors and $k/2$ edges connect to its leftmost neighbors.
  2. (2)
    To create $G_{t}$, for each node $[n]$, we take its $k/2$ rightmost neighbors and rewire them with probability $\beta$. For each of the $k/2$ rightmost neighbors that are to be rewired, we make one query to the LLM, which indicates how the edge will be rewired. The choice is made by providing the LLM with all the network nodes and each node’s neighbors (i.e., the network structure).
The model closely resembles the Watts-Strogatz model, with the primary distinction being the method of edge rewiring. Instead of randomly selecting edges for rewiring, as in the Watts-Strogatz model, we determine the rewiring of an edge by inquiring about the LLM and providing it with the current network structure.
We generate networks with varying sizes ($n=10$ to $n=100$) and examine the relationships between the average shortest path length ($L$) and the average clustering coefficient ($C$) with respect to the network size ($n$). We test $\beta\in\{0.25,0.5,0.75\}$ and set $k=5$ as an illustration. In addition, we compare with the results reported in (Jacobs and Watts, 2021). Some instances of these networks are shown in Figures 4(a), 4(b) and 4(c). We run regressions relating to the average shortest path length with $\log(n)$ and observe that for all temperatures, we obtain statistically significant results (see the significance indicators in Figure 5, where most regresssions have $P<0.001$) being that the average shortest path length scales proportionally to $\log(n)$. Similarly, we show that the average clustering coefficient scales as $1/\log(n)$ (again, most regression results have $P<0.001$). Both results follow the guideline from (Jacobs and Watts, 2021) which tested the small-world properties of organizational networks.
(a) $\beta=0.25$
(b) $\beta=0.5$
(c) $\beta=0.75$
Figure 5. Regression plots relating average shortest path length and average clustering coefficient with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$. $a$ represents the slope of the regression lines. *** indicates $P<0.05$, ***** * indicates $P<0.01$, and ******* * * indicates $P<0.001$.
However, a direct comparison with the Watts-Strogatz model reveals that the LLM-generated networks do not exactly mirror the statistics of Watts-Strogatz networks with corresponding $\beta$ values. As Figures 4(a), 4(b) and 4(c) show, we only observe weak evidence that the final generated networks have a similar average shortest path length as Watts-Strogatz with $\beta\in\{0.25,0.5,0.75\}$ ($P<0.1$, t-test comparing the resulting average shortest path lengths of the networks generated by multi-LLMs and the Watts-Strogatz average shortest path lengths for $\beta\in\{0.25,0.5,0.75\}$) and a larger average clustering coefficient than Watts-Strogatz ($P<0.1$, t-test comparing the resulting average clustering coefficient of the networks generated by multi-LLMs and the Watts-Strogatz average clustering coefficient for $\beta\in\{0.25,0.5,0.75\}$).
To quantify this resemblance, we conduct a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks. The results, depicted in Figures 4(d), 4(e) and 4(f), show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings. In Figures 4(d), 4(e) and 4(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$, each simulation and each temperature together with the $P$-values resulting from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$.
In conclusion, our analysis demonstrates that LLM-generated networks exhibit key small-world properties, with logarithmic scaling of average shortest path lengths and inverse logarithmic scaling of average clustering coefficients. While these networks do not perfectly align with the Watts-Strogatz model, they exhibit similar structural characteristics, underscoring the potential of LLMs in generating small-world networks.
2.3. Real-World Networks
To investigate the behavior of LLMs in real-world network formation contexts, we employ the Facebook100 dataset (Traud et al., 2012) and a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972). This dataset comprises “friendship” networks from one hundred American colleges and universities, captured at a specific moment in time from Facebook’s online social network. We model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives.
2.3.1. Discrete Choice Model
For each node $i_{t}$ that we consider at time $t$, we randomly remove one of its current friends from the real-world network. After we remove a neighbor for each of $i_{1},\dots,i_{T}$, we end up with the network $G_{1}$, which we use as a seed network for the LLM agents.
Subsequently, during the link formation process, we present each node $i_{t}$ with a set of candidate nodes (denoted by $A_{t}$), comprising one of the previously removed friends and other nodes that are not their friends. We then instruct the LLM to form a link with one of the candidates, providing the attributes of the candidates and the social network structure to aid its decision-making. These choices are made sequentially.
We use the utility of the model for each node for each sequential decision of network formation:
$U_{ij,t}=\theta_{\mathrm{PA}}\log d_{j,t}+\theta_{\mathrm{H}}\log w_{ij}+\theta_{\mathrm{TC}}\log c_{ij,t}+\epsilon_{ij,t}.$
In this equation, $\theta_{\mathrm{PA}}$ measures the strength of preferential attachment based on the degree $d_{j,t}$ of $j$ at step $t$, $\theta_{\mathrm{H}}$ measures the strength of homophily based on the similarity $w_{ij}$ (i.e. number of common attributes) between $i$ and $j$, and $\theta_{\mathrm{TC}}$ measures the strength of triadic closure, based on the number of common neighbors $c_{ij,t}$ between $i$ and $j$ at step $t$. The error term $\epsilon_{ij,t}$ is distributed as i.i.d. standard Gumbel [^9] The standard Gumbel distribution has CDF $e^{e^{-x}}$.. All variables are first normalized based on their range, and then the log transformation is taken.
The multinomial logit model (MNL) indicates that the probability that $i$ links to $j$ at step $t$ is given by
$p_{ij,t}=\Pr\left[\mathrm{argmax}_{r\in A_{t}}U_{ir,t}=j\right]=\frac{d_{j,t}^{\theta_{\mathrm{PA}}}w_{ij}^{\theta_{\mathrm{H}}}c_{ij,t}^{\theta_{\mathrm{TC}}}}{\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{ir}^{\theta_{\mathrm{H}}}c_{ir,t}^{\theta_{\mathrm{TC}}}}.$
Given a sequence of nodes $i_{1},\dots,i_{T}\in V$ and choices (denoted by subscripted $j$) $j_{1}\in A_{1},\dots,j_{T}\in A_{T}$, the parameters can be found by maximizing the log-likelihood function. To get the standard errors of the coefficients and the corresponding $P$-values, we follow the process outlined in (Overgoor et al., 2019).
2.3.2. Experiment Setup and Rationale
We use datasets that refer to three US colleges – Caltech36, Swarthmore42, and UChicago30. They represent different sizes of networks: $n=769$, $n=1,659$, and $n=6,591$ nodes, respectively.
For the Caltech36 ($n=769$) and Swarthmore42 ($n=1,659$) datasets, we consider all of the nodes of the networks as nodes that have to make a choice. For the UChicago30 dataset, we consider a randomly sampled subset of 2,000 nodes because of the limited context window of GPT models. We set the size of the alternative sets $A_{t}$ to 20, and thus, the accuracy of random guessing is 5%.
Figure 6. Multinomial logit coefficients for three networks from the Facebook100 dataset. The standard errors of the estimates are shown with error bars. The full regression tables can be found in Appendix A. All coefficients are statistically significant, where the null hypothesis corresponds to the respective parameter being equal to 0.
2.3.3. Results
We present the regression results in Figure 6101010More detailed results can be found in Table S.1 in Appendix A. Our analysis reveals that the degree, the number of common neighbors, and the number of shared interests are all statistically significant factors influencing the choices of LLMs ($P<0.001$) across all datasets. Furthermore, we find that LLMs place the greatest emphasis on the number of common attributes ($\hat{\theta}_{\mathrm{H}}$), followed by the number of common neighbors ($\hat{\theta}_{\mathrm{TC}}$), and then the degree ($\hat{\theta}_{\mathrm{PA}}$).
These results suggest that LLMs exhibit a strong preference for homophily, followed by a tendency towards triadic closure, and finally, preferential attachment. Specifically, the relationship $\hat{\theta}_{\mathrm{H}}>\hat{\theta}_{\mathrm{TC}}>\hat{\theta}_{\mathrm{PA}}$ holds across all examined datasets and temperature settings.
In addition to the coefficient $\hat{\theta}_{\mathrm{TC}}$ of triadic closure, we measure the percentage change in the average shortest path length and the average clustering coefficient between the network before the edge deletions and the network after the choices made by the LLM agents (see Table S.1). We observe that the average clustering coefficient also decreases compared to the initial network, indicating the weakening of triadic closure compared to the initial network.
We also provide reasoning analysis. Figure S.8 demonstrates that the rationales provided by LLMs aligns with the aforementioned findings. In particular, LLMs emphasize homophily, indicating that nodes sharing similar characteristics—such as student/faculty status, high school, and graduation year—are more likely to form connections. Conversely, the effects of triadic closure and preferential attachment are less prominent.
We also compare LLMs’ prediction with random guessing under a link prediction framework. We evaluate the percentage increase in accuracy, where accuracy is defined as the likelihood of selecting a previously existing neighbor from a set of alternatives, compared to making a random selection from 20 possible options. As shown in Appendix A, LLMs demonstrate a substantially higher accuracy in their selections than would be expected from random guessing.
Moreover, we examine whether the community structure is reinforced after the network formation decision made by LLMs. We run the Louvain algorithm 10 times for both the initial graph $G_{1}$ and the final graph $G_{T}$ and perform a t-test between the resulting modularities (see Appendix A). We observe that the modularity increases in all cases and is statistically significant. This suggests that LLMs make choices that naturally yield a reinforcement of the community structure.
To reflect the relation with “small-world” phenomenon, we measure the percent change in the average shortest path length between the network before the edge deletions and the network after the choices made by the LLM agents (see Appendix A). We observe a small decrease in the average path length in the new network in almost all cases indicating that the average shortest path length decreases slightly, indicating strengthening of the small-world phenomenon.
Taken together, our analysis indicates that although preferential attachment, homophily, and triadic closure are all present in the network formation behaviors of LLMs, homophily emerges as the predominant driving force. Moreover, both community structure and small-world phenomenon are reinforced.
3. Discussion
Our study conducted a comprehensive evaluation of LLMs’ network formation preferences, encompassing both micro-level network principles such as preferential attachment, triadic closure, and homophily and macro-level network principles like community structure and the small-world phenomenon. Our findings indicate that networks generated by multiple LLMs do exhibit these properties, particularly when primed with network statistics such as the number of mutual friends or the degree of alternatives. Furthermore, through discrete choice modeling, we explored the emergence of these properties in real-world networks. Our results reveal that the agents’ selections are predominantly driven by homophily, followed by triadic closure, and then preferential attachment. Additionally, we observed a strengthening of community structure in real-world network experiments.
Our study enhances our understanding of the behavior of multi-LLMs. For example, our findings reveal different strengths of network formation properties, suggesting that when LLMs are employed to coordinate social networks in social or work environments, they may exhibit behaviors similar to humans. This insight contributes to developing socially aware LLMs that align with human preferences, with potential applications ranging from chatbots and virtual assistants to collaborative systems and social media platforms.
Furthermore, our study may pave the way for innovative applications of LLMs in network science research. Given that LLMs adhere to fundamental network principles and exhibit behaviors akin to human network formation, they hold significant potential for use in simulations of social systems and agent-based modeling. LLM agents have the capability to produce more sophisticated and nuanced behaviors, including negotiation, collaboration, and competition. This opens up opportunities to simulate networked environments such as marketplaces or organizational structures, where LLMs can model decision-making processes and interaction patterns.
Looking ahead, we propose three promising directions for future research. First, we propose investigating LLM behavior in more complex interactions, such as simulated dialogs or specific environments like companies, schools, or parties. By examining LLMs in these contexts, we can better understand their network formation preferences and how they adapt to different social dynamics. This research could shed light on the nuances of LLM behavior in realistic settings and inform their integration into diverse social environments. Second, we suggest exploring the application of our findings in specific settings, particularly in organizational contexts. For example, considering the potential use of LLMs to assist HR professionals, our study’s insights could guide the alignment of LLMs with organizational goals. By leveraging network information, LLMs could help HR departments identify their firms’ best fit or talent, optimizing recruitment and talent management processes. Finally, we propose using our methods to create realistic synthetic networks. These synthetic datasets can serve as benchmarks for evaluating graph learning methods, thus addressing the scarcity of existing graph benchmarks. By adjusting parameters like temperature, we can generate diverse networks to test graph neural network performance under various conditions. In addition, our approach can generate artificial data resembling real-world data while adhering to privacy regulations. This has the potential to expedite research in graph learning and trustworthy machine learning by providing realistic network datasets generated by LLMs.
4. Methods
4.1. Experimental Procedure
In our study, we performed experiments to assess whether key network principles at both the micro-level (such as preferential attachment, triadic closure, and homophily) and the macro-level (including community structure and weak ties) align with classical network models. Subsequently, we utilized real-world networks to determine the factors that are most heavily weighted by LLMs.
Our experiments span a time series of $T$ steps, with a sequence of network structures denoted as $G_{1},G_{2},\dots,G_{T}$. The initial network, $G_{1}$, is referred to as the seed network. At each step $t$, we select a query node $i_{t}$ (which may either be a new arrival or an existing node in the graph) and assign it the task of forming new links. This is accomplished by selecting nodes from a set of alternatives $A_{t}$ (meaning potential candidates for link formation) and initiating a query call $\mathcal{Q}(A_{t},i_{t},\delta)$ to the LLM (as outlined in Algorithm 1) to create up to $\delta$ new links. The edge set selection process involves presenting the LLMs with personal or network features of the alternatives, denoted as $F(A_{t})=\left\{f_{a}:a\in A_{t}\right\}$, which may include information such as the neighbors of the nodes, node degrees, common connections with $i_{t}$, and community memberships, formatted in JSON. We adopt a zero-shot learning approach, avoiding the provision of examples to the model to prevent bias, in line with relevant studies such as (Brookins and DeBacker, 2023). This approach allows for the exploration of the innate preferences of LLMs.
We employ multiple temperatures to account for the variability in response generation by LLM systems, which is also observed in classical statistical models of network formation (Jackson, 2008). Our study conducts experiments using three temperatures: 0.5, 1.0, and 1.5.
Moreover, the model is tasked with outputting a JSON object indicating the node chosen for link formation and the rationale behind the choice. This approach is adopted because LLMs have demonstrated proficiency in processing code-like structures, such as HTML and JSON.
4.2. Summarizing the Reasoning of the LLM Agents
In addition to analyzing the innate preferences of LLM agents, we employ an interview-like approach to probe the rationale behind their decisions. Following the LLMs’ decisions on network formation, we inquire about their reasoning for making those choices. We then aggregate all the provided rationales using GPT models (GPT-3.5 for Section 2.1 and GPT-4 for Section 2.3) by asking it to identify the common themes, resulting in the creation of multiple categories. To determine the category to which each LLM’s reasoning belongs, we subsequently ask the GPT models to classify each rationale into the appropriate category.
4.3. Estimating the Parameters of the Discrete Choice Model
To estimate the parameters of the discrete choice model, we optimize the following log-likelihood function
$(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})=\operatorname*{arg\,max}_{(\theta_{\mathrm{PA}},\theta_{\mathrm{TC}},\theta_{\mathrm{H}})\in\mathbf{R}^{3}}\sum_{t=1}^{T}\left(\theta_{\mathrm{PA}}\log d_{j_{t},t}+\theta_{\mathrm{H}}\log w_{i_{t}j_{t}}+\theta_{\mathrm{TC}}\log c_{i_{t}j_{t},t}-\log\left(\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{i_{t}r}^{\theta_{\mathrm{H}}}c_{i_{t}r,t}^{\theta_{\mathrm{TC}}}\right)\right),$
where $i_{1},\dots,i_{T}$ are the chooser nodes (i.e., the LLM agents who want to form a link), and $j_{1},\dots,j_{T}$ are the nodes which are chosen from the alternative sets $A_{1},\dots,A_{T}$. The likelihood function is convex, and we optimize it with the L-BFGS-B method (Liu and Nocedal, 1989). The standard errors of the coefficients are approximated as $\sqrt{-H^{-1}}$ where $H$ is the Hessian matrix of the log-likelihood at $(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})$ (cf. (Overgoor et al., 2019; Train, 2009)).
Acknowledgements.
M.P. is supported by a scholarship from the Onassis Foundation (Scholarship ID: F ZT 056-1/2023-2024).
Appendix A Regression Table
In Table S.1, we report the regression coefficient for the regression in the real-world network data. The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 2.3 (also shown in Figure 6) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0). Next, LL corresponds to the log-likelihood of the fitted model, and AIC corresponds to the Akaike Information Criterion. Finally, we report the percent change in the accuracy compared to random guessing, the percent change in the average path length (as a measure of the small-world phenomenon), and the clustering coefficient (as a measure of the small-world phenomenon and the triadic closure), as well as the $t$-statistic for the change in modularity ($Q$) between the ground truth network dataset (before the edge deletions) and the network after the network formation process.
We observe that $\hat{\theta}_{\mathrm{H}}>\hat{\theta}_{\mathrm{TC}}>\hat{\theta}_{\mathrm{PA}}>0$ accross all settings. LLM agents do better than random guessing, reinforce the small-world phenomenon, and weaken the triadic closure. Finally, the community structure is strengthened after new links are formed.
Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ LL AIC % Change % Change % Change $\Delta Q$ (t-stat)
Acc. $L$ $C$
Caltech36 (769 nodes, 33,312 edges)
0.5 0.41*** (0.31) 1.95*** (0.64) 0.59*** (0.37) -1,377.47 2,762.94 171.8 -0.008 -9.94 3.45**
1.0 0.36*** (0.13) 1.85*** (0.52) 0.58*** (0.33) -1,435.07 2,878.13 179.6 -0.18 -11.08 3.49**
1.5 0.36*** (0.16) 1.72*** (0.31) 0.55*** (0.20) -1,522.47 3,052.94 127.6 -0.06 -11.46 3.37**
Swarthmore42 (1,659 nodes, 12,2100 edges)
0.5 0.18*** (0.13) 1.62*** (0.23) 0.65*** (0.08) -2,838.33 5,684.66 124.2 0.01 -11.46 7.42***
1.0 0.26*** (0.07) 1.70*** (0.35) 0.58*** (0.12) -2,927.99 5,863.97 91.6 -0.10 -4.25 1.96*
1.5 0.19*** (0.14) 1.50*** (0.33) 0.59*** (0.10) -3,139.42 6,286.83 87.39 -0.20 -4.52 4.03***
UChicago30 (6,591 nodes, 416,206 edges)
0.5 0.23*** (0.06) 2.00*** (0.24) 0.41*** (0.09) -3,444.33 6,896.67 217.2 -0.24 -2.52 7.46*** [0.34]
1.0 0.23*** (0.10) 1.98*** (0.17) 0.38*** (0.06) -3,578.18 7,164.36 219.2 -0.12 -2.66 9.56*** [1.05]
1.5 0.22*** (0.12) 1.78*** (0.27) 0.41*** (0.07) -2,033.49 4,074.98 222.4 -0.17 -2.42 10.19*** [0.24]
Notes $\hat{\theta}_{\mathrm{PA}}$ = Coefficient of log degree, $\hat{\theta}_{\mathrm{H}}$ = Coefficient of log # of common attributes, $\hat{\theta}_{\mathrm{TC}}$ = Coefficient of log # common neighbors
LL = Log-likelihood, AIC = Akaike Information Criterion
Acc. = Accuracy, $L$ = Average Path Length, $C$ = Average Clustering Coefficient, $\Delta Q$ (t-stat) = Modularity change t-statistic
:*${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
Table S.1. Multinomial logit coefficients for three networks from the Facebook100 dataset. The standard errors of the estimates are shown in parentheses. The null hypothesis corresponds to the respective parameter being equal to 0. We report the percent change in accuracy, average path length, and average clustering coefficient compared to the initial network (before the deletion of edges). For the change in modularity, we run the Louvain algorithm ten times and perform a t-test with the resulting modularities. For the UChicago30 dataset, we report the t-statistic value in the subgraph induced by the 2,000 sampled nodes, since the newly added edges would have a very small effect on the change in the community structure if we were to measure it in the whole network. We also report the modularity change (t-statistic) of the whole graph inside brackets.
Appendix B Analytical Regression Tables
We present the regression tables for all the combinations of coefficients for each of the three real-world network datasets. For all datasets, we observe that $\hat{\theta}_{\mathrm{PA}}$ is smaller than both $\hat{\theta}_{\mathrm{TC}}$ and $\hat{\theta}_{\mathrm{H}}$ in all models where any pair is included. Similarly, $\hat{\theta}_{\mathrm{TC}}$ is always smaller than $\hat{\theta}_{\mathrm{H}}$ in all models that are both included. Finally, note that whenever only $\hat{\theta}_{\mathrm{PA}}$ and $\hat{\theta}_{\mathrm{TC}}$ are considered, then $\hat{\theta}_{\mathrm{PA}}<0$ and the result is not statistically significant ($P>0.05$).
Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ Log Likelihood AIC
0.5 0.50*** (0.04) -2,236.36 4,476.71
0.5 2.78*** (0.07) -1,511.42 3,026.85
0.5 1.53*** (0.04) -1,506.01 3,016.02
0.5 0.64*** (0.06) 2.99*** (0.19) -1,414.71 2,835.41
0.5 -0.02 (0.10) 1.53*** (0.08) -1,505.95 3,017.90
0.5 1.43*** (0.11) 0.82*** (0.06) -1,406.39 2,818.78
0.5 0.41*** (0.31) 1.95*** (0.64) 0.59*** (0.37) -1,377.47 2,762.94
1.0 0.48*** (0.04) -2,242.85 4,489.70
1.0 2.69*** (0.07) -1,558.67 3,121.34
1.0 1.47*** (0.04) -1,556.68 3,117.37
1.0 0.58*** (0.05) 2.86*** (0.08) -1,473.13 2,952.26
1.0 -0.04 (0.06) 1.47*** (0.06) -1,556.24 3,118.48
1.0 1.40*** (0.28) 0.79*** (0.06) -1,457.76 2,921.52
1.0 0.36*** (0.13) 1.85*** (0.52) 0.58*** (0.33) -1,435.07 2,878.13
1.5 0.50*** (0.04) -2,233.25 4,470.50
1.5 2.51*** (0.07) -1,646.83 3,297.65
1.5 1.36*** (0.04) -1,636.20 3,276.40
1.5 0.57*** (0.07) 2.65*** (0.13) -1,559.57 3,125.15
1.5 -0.01 (0.05) 1.37*** (0.04) -1,636.19 3,278.37
1.5 1.29*** (0.11) 0.75*** (0.06) -1,546.78 3,099.55
1.5 0.36*** (0.16) 1.72*** (0.31) 0.55*** (0.20) -1,522.47 3,052.94
Note :*${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
Table S.2. Multinomial logit coefficients for Caltech36. The standard errors of the estimates are shown in parentheses.
Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ Log Likelihood AIC
0.5 0.33*** (0.03) -4,978.28 9,960.56
0.5 2.91*** (0.05) -3,027.03 6,058.06
0.5 1.37*** (0.02) -3,014.28 6,032.57
0.5 0.44*** (0.15) 2.99*** (0.10) -2,948.91 5,903.82
0.5 -0.18*** (0.09) 1.36*** (0.07) -3,002.17 6,010.35
0.5 1.42*** (0.11) 0.72*** (0.07) -2,847.16 5,700.32
0.5 0.18*** (0.13) 1.62*** (0.23) 0.65*** (0.08) -2,838.33 5,684.66
1.0 0.38*** (0.03) -4,959.02 9,922.04
1.0 2.83*** (0.05) -3,119.06 6,242.11
1.0 1.32*** (0.02) -3,118.13 6,240.26
1.0 0.50*** (0.16) 2.92*** (0.09) -3,018.85 6,043.71
1.0 -0.11** (0.09) 1.32*** (0.07) -3,113.21 6,232.43
1.0 1.41*** (0.15) 0.69*** (0.11) -2,947.89 5,901.78
1.0 0.26*** (0.07) 1.70*** (0.35) 0.58*** (0.12) -2,927.99 5,863.97
1.5 0.36*** (0.03) -4,952.51 9,909.02
1.5 2.64*** (0.05) -3,324.71 6,653.43
1.5 1.24*** (0.02) -3,306.06 6,616.11
1.5 0.44*** (0.11) 2.71*** (0.05) -3,241.67 6,489.35
1.5 -0.14*** (0.10) 1.24*** (0.05) -3,298.07 6,602.13
1.5 1.30*** (0.14) 0.67*** (0.09) -3,150.29 6,306.57
1.5 0.19*** (0.14) 1.50*** (0.33) 0.59*** (0.10) -3,139.42 6,286.83
Note :*${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
Table S.3. Multinomial logit coefficients for Swarthmore42. The standard errors of the estimates are shown in parentheses.
Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ Log Likelihood AIC
0.5 0.28*** (0.02) -5,983.56 11,971.13
0.5 2.93*** (0.05) -3,637.18 7,278.37
0.5 1.11*** (0.02) -3,740.78 7,485.55
0.5 0.38*** (0.04) 3.06*** (0.21) -3,523.06 7,052.12
0.5 -0.04 (0.08) 1.11*** (0.04) -3,739.31 7,484.61
0.5 1.68*** (0.13) 0.51*** (0.07) -3,477.37 6,960.75
0.5 0.23*** (0.06) 2.00*** (0.24) 0.41*** (0.09) -3,444.33 6,896.67
1.0 0.28*** (0.02) -5,982.66 11,969.32
1.0 2.85*** (0.05) -3,759.78 7,523.56
1.0 1.07*** (0.02) -3,879.04 7,762.08
1.0 0.36*** (0.07) 2.96*** (0.31) -3,649.84 7,305.68
1.0 -0.04 (0.06) 1.07*** (0.03) -3,877.83 7,761.66
1.0 1.67*** (0.11) 0.49*** (0.08) -3,611.48 7,228.95
1.0 0.23*** (0.10) 1.98*** (0.17) 0.38*** (0.06) -3,578.18 7,164.36
1.5 0.30*** (0.03) -3,241.67 6,487.34
1.5 2.71*** (0.06) -2,145.02 4,294.03
1.5 1.03*** (0.02) -2,175.32 4,354.64
1.5 0.37*** (0.08) 2.81*** (0.10) -2,080.67 4,167.34
1.5 -0.01 (0.04) 1.03*** (0.04) -2,175.25 4,356.49
1.5 1.50*** (0.10) 0.51*** (0.04) -2,051.61 4,109.23
1.5 0.22*** (0.12) 1.78*** (0.27) 0.41*** (0.07) -2,033.49 4,074.98
Note :*${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
Table S.4. Multinomial logit coefficients for UChicago30. The standard errors of the estimates are shown in parentheses.
Appendix C LLM Reasoning Summarization
In addition to studying LLM agents’ innate preferences, we use an interview-style approach to understand why they make certain decisions. After LLMs form networks, we ask them about their reasoning. We then use GPT models (GPT-3.5 for Section 2.1 and GPT-4 for Section 2.3) to group their explanations into common themes, creating various categories. To classify each LLM’s reasoning, we have the GPT models assign them to the relevant category.
C.1. Micro-level Principles
In Figure S.7, we show the reasoning of the LLM agents from the networks of Section 2.1.
In Figures 7(a) and 7(b), the reasons mean the following:
  • Well-Connectedness: it corresponds to reasons related to the person being well-connected and having many friends.
  • Mutual Friends: it is related to having many friends in common with the person so that the person can introduce them to their friends.
We observe that for Figure 7(a), the well-connectedness reason becomes less important as the temperature increases while having more mutual friends becomes more important as the temperature increases. This means that the agents emphasize triadic closure in higher temperatures and decrease emphasis on preferential attachment. For Figure 7(b), we observe an opposite pattern in the LLM reasoning. A possible explanation for this asymmetry is the fact that different types of features are provided, i.e., in the case of Figure 7(a) the neighbors of each node are provided, whereas in the case of Figure 7(b) the common neighbors with the query node $i_{t}$ are provided.
Finally, Figure 7(c) contains the reasons for Figure 3(a) based on having a common hobby, a common color, or a common location. We observe that the LLM agents make choices primarily based on sharing the same hobby and then based on a common location or a common color (on an almost equal basis).
(a) Reasons for Figure 1(a)
(b) Reasons for Figure 2(a)
(c) Reasons for Figure 3(b)
Figure S.7. Distribution of reasoning across categories for the networks of Figures 1(a), 2(a) and 3(b). The classification uses GPT-3.5 on 20 randomly sampled reasons repeated over five times (micro-averaged frequencies). The error bars correspond to 95% confidence intervals.
C.2. Real-world Data
We also present the LLM reasoning for real-world networks. The reasons have been set so that there is a one-to-one correspondence with the regression coefficients. Specifically, the Mutual Friends reason corresponds to $\hat{\theta}_{\mathrm{TC}}$, the Number of friends reason to $\hat{\theta}_{\mathrm{PA}}$, and the Similar attributes reason corresponds to $\hat{\theta}_{\mathrm{H}}$ respectively. Consistently with the results reported in Table S.1, the agents’ reasoning is more geared towards homophily (same attributes), then triadic closure (mutual friends), and finally preferential attachment (number of friends). These results are consistent with the regression.
Figure S.8. Distribution of reasoning across categories for the real-world networks. The classification is done using GPT-4 on a set of 20 randomly sampled reasons repeated over 5 times. In agreement with Table S.1, the multi-LLMs put more weight on homophily, then on triadic closure, and then on preferential attachment. The error bars correspond to 95% confidence intervals.
Appendix D Network Evolution
Here we depict the evolution of the networks generated by the LLM agents.
D.1. Principle 1: Preferential Attachment
We plot the evolution of the LLM-based preferential attachment networks at three timesteps, together with the degree distribution alongside the degree distribution of a BA graph with the same number of nodes. We observe that for the temperature being 0.5 we have a core-periphery like formation which diverges from the BA model, whereas for the temperature being 1.5 the network has the same degree distribution as the BA model.
Figure S.9. Dynamic evolution of networks created based on Principle 1.
D.2. Principle 2: Triadic Closure
We plot the evolution of the LLM-generated networks based on the triadic closure principle, together with the transitivity measure and the algebraic connectivity (which corresponds to the second-smallest eigenvalue of the graph Laplacian). We observe that the algebraic connectivity gradually increases as new edges between the clusters are created. Specifically, the algebraic connectivity reaches a higher value for higher temperatures, indicating the more frequent creation of new intra-cluster edges. Moreover, we observe that the transitivity initially increases and then decreases until it reaches its final value.
Figure S.10. Dynamic evolution of networks created based on Principle 2.
Appendix E Feature Representations for Prompts
Below, we give examples of the features used in the prompt presented in Algorithm 1. The features are formatted as a list of JSON objects which are provided to the prompt.
E.1. Principle 1: Preferential Attachment
For the structure-based information (Figure 1(a)) we have the following features:
            "name" : 0,
            "neighbors" : [5, 7, 1, 6]
And for the degree-based information (Figure 1(b)) we have the following features:
            "name" : 0,
            "degree" : 4
E.2. Principle 2: Triadic Closure
For the structure-based information (Figure 2(a)) we have the following features:
            "name" : 0,
            "common_neighbors" : [5, 7, 1, 6]
And for the degree-based information (Figure 2(b)) we have the following features:
            "name" : 0,
            "num_common_neighbors" : 4
E.3. Principle 3: Homophily
We have the following features:
            "name" : 0,
            "favorite_color" : "red",
            "hobby" : "hiking",
            "location" : "Boston"
E.4. Principle 5: Small-World
We have the following features:
            "name" : 0,
            "neighbors" : [5, 7, 1, 6]
E.5. Real-World Data
We have the following features:
            "name" : 0,
            "status" : "student",
            "major" : 10,
            "second major" : 93,
            "accommodation" : "house",
            "high_school" : 5,
            "graduation_year" : 2008
We note that the initial Facebook100 dataset included gender information as a feature. We chose not to include the gender as one of the features as it has been shown that language models exhibit gender bias (Vig et al., 2020; Kotek et al., 2023; Bordia and Bowman, 2019).

### v3

@@
 We also investigate LLMs’ decision-making based on real-world networks, revealing that triadic closure and homophily have a stronger influence than preferential attachment and that LLMs perform well in network formation predictions.
-Overall, our study opens up new possibilities for using LLMs in network science research and helps develop socially aware LLMs by shedding light on their network formation behaviors and exploring their impacts on social dynamics.
+Overall, our study opens up new possibilities for using LLMs in network science research and helps develop socially aware LLMs by shedding light on their social interaction behaviors and exploring their impacts on social dynamics.
 large language models, networks, network formation, link prediction, multi-agent systems
 conference: ; ;
-1. Introduction
+Introduction
 Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023) and Llama 2 (Touvron et al., 2023), have shown promising developments in AI techniques and their integration into real-life applications.
@@
 Misaligned AI actions may lead to unintended consequences, such as biased decision-making, fairness issues, and the miscoordinative or non-cooperative behavior (Rahwan et al., 2019).
-Recently, researchers have started to apply methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods, to study LLMs.
-These methods help uncover the generative capabilities of LLMs and their interpretability (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
+Recently, researchers have started to apply social science methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods, to study LLMs.
+These methods not only reveal the capabilities and interpretability of LLMs but also suggest their potential for applications in social science (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
 In human societies, social networks play a crucial role in shaping individual behaviors, preferences, and connections, as well as influencing the diffusion of information and norms across communities (Rogers, 2003; Bakshy et al., 2012; Fowler and Christakis, 2008; Banerjee et al., 2013; Yuan et al., 2018).
-LLMs have shown great potential in social contexts, notably as intelligent personal assistants facilitating social and business interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)).
+LLMs have shown great potential in social contexts, notably as intelligent personal assistants that facilitate social and prosocial interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)).
 However, less is known about how LLMs’ innate behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023).
-This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the intricate networks of human relationships, which is a fundamental aspect of social systems.
-Our study thus explores LLMs’ behaviors and preferences in network formation in both synthetic and real-world social networks.
-By analyzing interactions between multiple LLMs (or multi-LLMs), we aim to understand the implications of LLMs representing humans in social and business settings.
+This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the networks of human relationships, which is a fundamental aspect of social systems.
+Our study thus explores LLMs’ behaviors and preferences in the context of network formation with both synthetic and real-world social networks.
+By analyzing interactions between multiple LLMs (or multi-LLMs), we aim to understand the implications of LLMs representing humans in social and professional settings.
 Specifically, we examine micro-level social network properties including preferential attachment (Barabási and Albert, 1999), triadic closure (Granovetter, 1973), and homophily (McPherson et al., 2001), as well as macro-level properties including community structure (Newman and Girvan, 2004), and the small-world phenomenon (Kleinberg, 2000; Watts and Strogatz, 1998).
-In synthetic networks, our findings reveal that LLMs can generate networks exhibiting preferential attachment and following scale-free degree distributions.
+In simulations conducted using synthetic networks, we have the following findings: First, we observe that LLMs can generate networks exhibiting preferential attachment and following scale-free degree distributions.
 Moreover, LLMs demonstrate strong tendencies toward triadic closure and exhibit tendency of homophily.
-These tendencies further explain the emergence of community structures with high modularity in networks formed by multiple LLMs.
+These tendencies further explain the emergence of community structures in networks formed by multi-LLMs.
 Additionally, we show that LLM agents exhibit dynamics consistent with small-world phenomena akin to those observed in the Watts-Strogatz model (Watts and Strogatz, 1998).
-Next, we conduct a comparative analysis of these principles to identify the key factors driving network dynamics in the context of real-world network formation scenarios.
-To do so, we consider real-world networks and ask each LLM to simulate a node in the network.
+Next, we compare between these principles to identify the key factors driving network dynamics in real-world network formation contexts.
+To achieve this, we consider real-world networks and ask each LLM to simulate a node in the network.
 We provide information on their network structure and ask for their decisions for the next link formation or dissolution step.
@@
 Additionally, the resulting choices of LLM agents strengthen the network’s community structure.
-Our study contributes to the literature on LLMs in several ways.
-First, it advances the discussion on aligning LLMs with human social preferences, which has potential applications in developing socially aware LLMs for real-life interactions.
+Our study contributes to the literature on LLMs in multiple ways.
+First, it advances the discussion on aligning LLMs with human preferences, which has potential applications in developing socially aware LLMs for real-life interactions.
 Insights into LLMs’ innate behaviors in network formation can inform the design of LLMs for effective collaboration in social and professional settings.
-Second, our work has direct applications to synthetic dataset generation, a problem important for the graph learning community.
+Second, our study could contribute to understanding the impact of LLMs on social dynamics and norms, such as their role in shaping online communities’ interaction patterns or influencing the adoption of new communication norms in professional settings.
+Finally, our work has direct applications to synthetic dataset generation, a problem important for the network science community.
 By controlling for a few parameters, such as the temperature, our work can be utilized to generate a variety of datasets, as well as generate synthetic datasets from real-world data to protect privacy.
-Finally, our study could contribute to understanding the impact of LLMs on social dynamics and norms, such as their role in shaping online communities’ interaction patterns or influencing the adoption of new communication norms in professional settings.
-    # Task     Your task is to select a set of people to be friends with.
-# Profile     Your profile is given below after chevrons:     <PROFILE>$F(\{i_{t}\})$</PROFILE>     # Candidate Profiles     The candidate profiles to be friends with are given below after chevrons:     <PROFILES>$F(A_{t})$</PROFILES>     # Output     The output should be given a list of JSON objects with the following structure     [         {{             "name" : name of the person you selected,             "reason" : reason for selecting the person         }}, ...
-]     # Notes     - The output must be a list of JSON objects ranked in the order of preference.
-- You can make at most $\delta$ selections.
-ALGORITHM 1 Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$ formatted as a Python f-string.
 (a) Structure-based
@@
 Figure 1.
-Networks with $n=200$ nodes created based on Principle 1 for various temperatures.
-For the structure-based networks (Figure 1(a)), we provide the multi-LLMs with the neighborhoods $\{N_{j,t}:j\in i_{t}\}$, and for the degree-based networks (Figure 1(b)), we provide the degrees $\{d_{j,t}:j\in i_{t}\}$ of the nodes $j\in i_{t}$.
-We also plot the degree distribution and compare it against a BA network with $n=200$ nodes.
-Top: In Figure 1(a), we observe that the resulting networks follow a power-law degree distribution ($P>0.5$, Bootstrap hypothesis test).
-When the temperature is 1.5, the resulting network resembles the BA model ($P<0.001$, K-S test).
-Bottom: When the degree information is provided only (Figure 1(b)), the resulting networks correspond to unrealistic star-like formations that are far from a scale-free network and resemble more a core-periphery network.
-2. Results
-Our study employs the state-of-the-art LLMs, such as GPT-3.5 and GPT-4 (OpenAI, 2023) with the OpenAI API [^1] For GPT-3.5 we use the gpt-3.5-turbo model provided by OpenAI, and for GPT-4 we use the gpt-4-1106-preview model..
-We constructed multiple LLMs using multiple independent conservation threads where each thread represents one multi-LLM.
-We prompt each LLM by providing information about other LLMs’ network information as well as decisions they are faced with.
-We test both state-of-the-art GPT-based models: for the results of Sections 2.2 and 2.1, we use GPT-3.5 agents, and in Section 2.3, we use GPT-4 agents.
-We obtained similar results using GPT-3.5 and GPT-4.
-Overall, we adopted a standard framework for multi-LLMs but underscored the multi-agent nature of the design in our setting (Leng et al., 2023; Akata et al., 2023; Brookins and DeBacker, 2023; Aher et al., 2023; De Marzo et al., 2023).
-We evaluate the robustness of our results by testing various temperature settings in LLMs, which control the degree of variation of the model output.
-All experiments employ a dynamic network setup.
-We also employ an interview-like approach to inquire about the rationale behind the decisions of LLMs.
-The details of our methods are presented in Methods.
-In this section, we first discuss three major micro-level network principles: preferential attachment, triadic closure, and homophily.
-We then examine two macro-level network principles: community structure and small-world phenomenon.
-Subsequently, we test the strength of these principles when LLM agents face real-world networks and compare the preferences in network formation tasks between LLMs and humans.
-2.1. Micro-Level Properties
+Results for Principle 1 (preferential attachment): We display simulated networks comprising 200 nodes across different temperatures.
+For structure-based simulations (Figure 1(a)), the multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$, while for degree-based simulations (Figure 1(b)), node degree data $\{d_{j,t}:j\in V_{t}\}$ was provided ($V_{t}$ corresponds to the vertex set of the network $G_{t}$ at round $t$).
+The degree distributions were then compared with those of a BA network with $n=200$ nodes.
+(a): Power-law degree distributions are evident ($P>0.5$, K-S test), with the networks at a temperature of 1.5 closely resembling the BA model ($P>0.1$, K-S test).
+(b): With degree information only, the networks form more unrealistic star-like structures, diverging from scale-free configurations and more closely mirroring a core-periphery network structure.
+Results
+In our study, we leveraged state-of-the-art LLMs through the OpenAI Application Programming Interface (API).
+Our methodology involved constructing multiple LLMs, each acting as an independent agent within separate conversational threads.
+Each LLM agent was prompted with network information about its peers and faced with decision-making scenarios.
+For synthetic network simulations, we employed GPT-3.5 agents, whereas GPT-4 agents were used for real-world network analysis.
+[^1] We utilized the gpt-3.5-turbo model for GPT-3.5 and the gpt-4-1106-preview model for GPT-4, both provided by OpenAI.
+Notably, both GPT-3.5 and GPT-4 yielded consistent results.
+We adopt a multi-LLM framework, emphasizing the multi-agent aspect of our experimental design, which shares similar spirits with recent research in the field (Leng et al., 2023; Akata et al., 2023; Brookins and DeBacker, 2023; Aher et al., 2023; De Marzo et al., 2023).
+To assess the robustness of our findings, we varied the ‘temperature’ settings in LLMs, which modulate the output variability, in a dynamic network environment.
+Additionally, we adopted an interview-like method to probe the rationale behind the LLMs’ decision-making processes, enriching our understanding of their behavior in network simulations.
+In our analysis, we begin with an examination of three fundamental micro-level network principles: preferential attachment, triadic closure, and homophily.
+Subsequently, we shift our focus to two macro-level network phenomena: community structure and the small-world phenomenon.
+Finally, we assess the resilience of these principles as LLM agents interact with real-world networks.
+This allows us to compare and contrast the network formation preferences between LLMs and humans.
+Micro-Level Properties
 Principle 1: Preferential Attachment
-Preferential attachment is a well-known mechanism in network science that explains how certain nodes in a network become more connected over time, leading to the emergence of highly connected vertices and scale-free distribution of degree (Barabási and Albert, 1999; Bianconi and Barabási, 2001).
-We examine whether LLMs exhibit the preferential attachment mechanism in the following way: First, we start with an empty network and add nodes at each time step $t$.
-For each new time step, we add a new node $i_{t}$, providing information on the network neighbors of all existing nodes.
-[^2] See Appendix E, for information about how the features are represented in the prompt.
-We chose to generate networks with $n=200$ nodes, as this size is sufficiently large to allow for the observation of degree distribution patterns.
-We plot one of the examples of resulting networks under each temperature in Figure 1(a).
-As observed in all temperature setups, we observe a few nodes that have many connections, and the majority of the notes have one or few connections, which renders the degree distribution scale-free, i.e., the probability $\pi(d)$ a node has degree $d$ is given by
-$\pi(d)\propto d^{-\gamma},\quad\text{where }\gamma>1.$
-Furthermore, we observe mode “super nodes” with many connections as the temperature increases.
-We also compare with the Barabási–Albert (BA) model (Barabási and Albert, 1999) where each incoming node $t$, the probability $p_{i,t}$ of it connecting to node $i$ is proportional to ${d_{i,t}}$, where $d_{i,t}$ is the degree of node $i$ at time step $t$, namely,
-$p_{i,t}=\frac{d_{i,t}}{\sum_{j=1}^{t}d_{j,t}}.$
-For temperatures of 0.5 and 1.0, we find that the estimated exponent values are $\hat{\gamma}=1.50\pm 0.23$ and $1.92\pm 0.24$, respectively.
-However, the degree distributions do not resemble the degree distribution of the generated BA network with an expected exponent $\gamma=3$.
-This is supported by the $P$-values from a two-sample Kolmogorov-Smirnov (K-S) test, which are $P<0.001$ and $P<0.05$, respectively when compared to the empirical degree distribution of the BA network [^3] As shown by (Barabási, 2013), the BA model follows a scale-free distribution of the form $\pi_{\mathrm{BA}}(d)\propto d^{-3}$ for sufficiently large $n$.
-However, because our $n$ is small, so that the prompt does not exceed the token limit for the GPT-3.5 model, we directly compared it to a sample from the BA model (the generated network has $\hat{\gamma}_{\mathrm{BA}}=2.24\pm 0.24$), instead of comparing with the limiting distribution $\pi_{\mathrm{BA}}(d)$..
-Moreover, for all temperatures, the resulting degree distribution is scale-free ($P>0.5$, bootstrapping hypothesis test comparing with the fitted scale-free distribution).
-We also examine the network formation when only degree information (rather than network structure) is provided, in a similar spirit of (De Marzo et al., 2023).
-We present the results in Figure 1(b).
-We observe “star-like” formations around one node.
-The degree distributions differ from scale-free distributions produced by statistical models such as the BA model ($P<0.001$, two-sample K-S test compared to the empirical distribution of the generated BA network).
-They are closer to a core-periphery network (Borgatti and Everett, 2000).
-Finally, we employ text analysis of reasoning provided by the agents for the results of Figure 1(a) [^4] We do not provide the distribution of reasons for the results of Figure 1(b) as all of the LLM’s decisions are made given the degree information..
-We observe two main categories of reasons (see Appendix C) for the fraction under each temperature):
-  • Well-Connectedness: it corresponds to reasons related to the person being well-connected and having many friends.
-  • Mutual Friends: it is related to having many friends in common with the person so that the person can introduce them to their friends.
-The well-connectedness reason becomes less prevalent as the temperature increases, and the mutual friends’ reason becomes more prevalent.
-Note that when the temperature is 1.5, the two reasons have almost the same frequency.
-Taken together, networks created by LLM systems, when given a network structure, demonstrate preferential attachment and exhibit scale-free degree distributions.
+Preferential attachment is a fundamental concept in network science, illustrating how nodes in a network gain connections over time, leading to a scale-free degree distribution characterized by a few highly connected nodes (Barabási and Albert, 1999; Bianconi and Barabási, 2001).
+In our study, we investigate whether LLMs manifest the preferential attachment mechanism.
+Starting with an empty network, we introduce nodes sequentially at each time step $t$.
+A new node $i_{t}$ is added during time step $t$, with the network’s existing nodes providing context about their connections.
+We generated networks with $n=200$ nodes, a size large enough to observe meaningful patterns in the degree distribution.
+Our simulations include two distinct types: structure-based, where the network’s structure was provided, and degree-based, where only the degree information of potential connections was given.
+[^2] For details on how features were represented in the prompts, refer to SI Section 5.
+The graphical outputs from our structure-based simulation, as illustrated in Figure 1(a), reveal a distinct pattern where a small number of nodes possess a high number of connections, whereas the majority exhibit fewer links.
+This pattern suggests a scale-free degree distribution, characterized by the probability $\pi(d)$ that a node has degree $d$, mathematically represented as:
+(1) $\pi(d)\propto d^{-\gamma},\quad\text{where }\gamma>1.$
+Interestingly, we also observe that the number of highly connected “super nodes” increases with temperature.
+[…]
+Furthermore, in degree-based simulations in a similar spirit of ((De Marzo et al., 2023); see Figure 1(b)), we noted formations resembling a “star” around central nodes.
+These distributions deviate from the scale-free patterns seen in statistical models like the BA model ($P<0.001$, two-sample K-S test), aligning more closely with core-periphery structures (Borgatti and Everett, 2000).
+To interpret the LLM agents’ decision-making, we analyzed the reasoning they provided, focusing on the structure-based results (Figure 1(a)) [^4] The degree-based results’ reasoning is not included as decisions were based solely on degree information..
+Two primary reasons emerged, varying with temperature (detailed in SI Section 3):
+  • Well-Connectedness: Relating to the node’s degree (number of connections).
+  • Mutual Friends: Relating to the number of shared connections.
+Interestingly, as the temperature increased, the prevalence of the well-connectedness rationale decreased, while the mutual friends’ rationale increased.
+In conclusion, our analysis demonstrates that LLMs, when provided with network structure, can replicate the preferential attachment phenomenon.
 Principle 2: Triadic Closure
@@
 Figure 2.
-Networks created based on Principle 2.
-Top: Figure 2(a) shows the resulting networks created by Principle 2 when the intersection of the neighborhoods of the query node and each alternative is provided.
-The node colors correspond to the stochastic block model (SBM) groups each node belongs to.
+Results for Principle 2 (triadic closure).
+(a) Figure 2(a) shows the resulting networks created by triadic closure when the intersection of the neighborhoods of the query node and each alternative is provided.
+The node colors correspond to the stochastic block model (SBM) groups to which each node belongs.
 The red edges correspond to the newly created edges between nodes of the red cluster, the blue edges correspond to the newly created edges between nodes of the blue cluster, and the orange edges correspond to edges created between the two clusters.
@@
 We observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test).
-Bottom: Figure 2(b) shows the same networks with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative.
+(b) Figure 2(b) shows the same networks with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative.
 Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test).
-The second micro-level principle we examine is triadic closure, which describes the phenomenon that individuals are likely to form connections with their friends’ friends, leading to closed triads within the network.
-This tendency enhances the network’s structure and cohesion, rooted in the likelihood of two nodes forming a direct connection if they share a common neighbor (Granovetter, 1973; Mosleh et al., 2024).
-To test for triadic closure, we use an assortative stochastic block model (SBM) (Newman, 2003) to generate an initial seed network $G_{1}$ with $n$ nodes divided into two equal-sized clusters $A$ and $B$.
-Connections within each cluster have a probability of $0.5$, while connections across clusters have a probability of $0.1$.
-This setup reflects our intuition that nodes within the same cluster are more likely to form connections due to their higher number of common neighbors.
-We then iterate over each node $i$, providing features based on the intersection of the neighborhoods of all of $i$’s non-neighbors or the number of common friends between $i$ and any non-neighbor $j$.
-This modeling choice contrasts the results mentioned about preferential attachment, where the neighborhood information is provided and made to make the two settings consistent [^5] We want to note that performing the same experiment when we provide the neighbors instead of the common neighbors produces similar results..
-We compare this with the random null model, which selects a node at random to connect to.
-We run ten simulations for $n=50$ nodes.
-[^6] For better visualization, we chose $n=50$ nodes instead of a large number of nodes such as $n=200$.
-Moreover, the results are statistically significant even with $n=50$ nodes.
-To measure whether the principle of triadic closure exists, we employ the fraction $\hat{p}$ of newly added edges between nodes of the same community, i.e., the fraction of edges in $G_{T}\setminus G_{1}$ that have their ends both at either $A$ or $B$, namely
+The error bars correspond to 95% confidence intervals.
+The second micro-level principle we examine is triadic closure, which posits that individuals are more likely to form connections with friends of friends, thus creating closed triads in the network.
+This process strengthens network structure and cohesion, grounded in the idea that two nodes are more likely to connect if they share a common neighbor (Granovetter, 1973; Mosleh et al., 2024).
+To investigate triadic closure, we employ an assortative stochastic block model (SBM) (Newman, 2003) to create an initial network $G_{1}$ with $n$ nodes divided into two equal-sized clusters $A$ and $B$.
+Connections within each cluster are formed with a probability of $0.5$, while inter-cluster connections occur with a probability of $0.1$.
+This setup mirrors our assumption that nodes within the same cluster are more inclined to connect due to a higher number of shared neighbors.
+In subsequent time steps, we then examine each node $i$, considering the intersection of neighborhoods of $i$’s non-neighbors or the number of mutual friends between $i$ and any non-neighbor $j$ [^5] It’s worth noting that similar outcomes arise when providing neighbors instead of common neighbors.
+We conducted ten simulations with $n=50$ nodes to facilitate clear visualization and ensure statistical significance [^6] Choosing $n=50$ instead of a larger number like $n=200$ aids in visualization and maintains statistical significance..
+To contrast LLM-based network formation, we refer to a random null model, selecting nodes at random for connection.
+For evaluating triadic closure, we utilize two metrics: marginal transitivity and probability of edge formation within the same community.
+Marginal transitivity ($D$) represents the change in the ratio of closed triangles to all triads, transitioning from the initial network $G_{T}$ to the SBM-generated $G_{1}$:
+$D=3\times\frac{\text{\# triangles}(G_{T})}{\text{\# triads}(G_{T})}-3\times\frac{\text{\# triangles}(G_{1})}{\text{\# triads}(G_{1})}.$
+where a large positive $D$ indicates a strong triadic closure tendency.
+The probability of forming an edge within the same community ($\hat{p}$) is calculated by the ratio of edges in $G_{T}\setminus G_{1}$ (newly formed edges) connecting nodes within the same cluster:
 $\hat{p}=\frac{\left|\left\{\{i,j\}\in E(G_{T})\setminus E(G_{1}):y_{i}=y_{j}\right\}\right|}{\left|E(G_{T})\setminus E(G_{1})\right|},$
-where $y_{i},y_{j}\in\{A,B\}$ are the community memberships of nodes $i$ and $j$ respectively, and $E(G_{T})\setminus E(G_{1})$ corresponds to the newly added edges.
-Generally, a value of $\hat{p}>0.5$ indicates the tendency of triadic closure since an LLM agent, which is supposed to exhibit triadic closure given the common neighbors, would ideally create a link within the same cluster.
-Additionally, we introduce the concept of marginal transitivity for the network, represented by $D$.
-This metric is defined as the change in the proportion of closed triangles to open triads in the network, from $G_{T}$ (the initial network) to $G_{1}$ (which corresponds to the SBM):
-$D=3\times\frac{\text{\# triangles}(G_{T})}{\text{\# triads}(G_{T})}-3\times\frac{\text{\# triangles}(G_{1})}{\text{\# triads}(G_{1})}.$
-We depict the sample outcomes in Figure 2(a), where the upper panel presents the results when the entire network structure is provided, whereas the lower panel only provides the number of common neighbors.
-Node colors correspond to their respective groups within the SBM.
-Red edges denote connections within the red cluster, blue edges within the blue cluster, and orange edges between the two clusters.
-Newly established edges are emphasized with a thicker line width.
-We present the probabilities of network formation within a community in the last panels in each row.
-We find that the probability of forming a within-community tie across all temperatures is substantially larger than the random network formation scenario ($P<0.001$, t-test comparing to Random), suggesting LLMs’ tendency of triadic closure (since an LLM which is given the common neighbors as features should ideally create links within the same community).
-Additionally, we observe a statistically significant increase in the marginal transitivity $D$ for all temperatures ($P<0.001$, t-test).
-Moreover, in all cases, the marginal transitivity $D$ is statistically significant compared to Random ($P<0.001$, t-test).
-These results suggest that LLM agents exhibit the tendency of triadic closure.
-Similar to the approach in earlier reasoning analysis, we present the most common reasons provided by LLMs in Appendix C [^7] We do not provide the distribution of reasons for the results of Figure 2(b) as all of the LLM’s decisions are made given the number of common neighbors information..
-We find that most of the agents base their decisions on the number of mutual friends.
+where $y_{i},y_{j}\in{A,B}$ denote the community memberships of nodes $i$ and $j$, respectively.
+A value of $\hat{p}$ exceeding $0.5$ suggests a triadic closure tendency.
+As we investigate under SBM, same community membership indicates more open triads being closed.
+In Figure 2(a), sample networks are displayed, with the upper panel showing networks where the entire structure is provided and the lower panel showing those with only common neighbor numbers provided.
+Nodes are color-coded to indicate their cluster memberships in the SBM, with red and blue edges within clusters and orange edges between clusters.
+Newly formed edges are highlighted with thicker lines.
+Marginal transitivity, presented in the fourth panel of each row, shows a statistically significant increase across all temperatures ($P<0.001$, t-test), and compared to the Random model, the increase is notable ($P<0.001$, t-test).
+Regarding the probability of within-community network formation (last panels in each row), we found that across all temperatures, this probability significantly exceeds that in the random network formation scenario ($P<0.001$, t-test against Random), indicating an LLM tendency towards triadic closure.
+These findings affirm that LLM agents tend to follow triadic closure.
+We investigate the reasons provided by LLMs for their decisions in SI Section 3.777Reasons for decisions based on the number of common neighbors are not provided, as these decisions are solely based on this information.
+The analysis revealed that many decisions were influenced by the number of mutual friends.
 Principle 3: Homophily
-(a) Assortativity Coefficients
-(b) Network Instances
-(c) Community Structure
-Figure 3.
-Top: Figure 3(a) shows the assortativity coefficients for each network attribute created based on Principle 3.
-Each node can form at most $\delta=5$ links with the other nodes based on their attributes.
-We observe that for all temperatures, we have positive assortativity ($P<0.001$, t-test with 0), which indicates homophily.
-The assortativities are significantly different for all temperatures from random or heterophilous ($P<0.001$, t-test comparing with Random) Middle: Networks created based on Principle 3.
-Each node can form at most $\delta=5$ links with the other nodes based on their attributes.
-The edge widths correspond to the number of common attributes shared between the nodes.
-We report the probability of forming a link as a function of the number of common attributes between two nodes.
-The node colors correspond to the communities found by running the Louvain algorithm.
-Bottom: Louvain modularity, number of communities, and community sizes of networks created based on Principle 3.
-Lower temperatures yield higher modularity ($P<0.001$, t-test) and fewer and larger communities ($P<0.001$, t-test).
-For smaller temperatures, the communities are more “well-separated”.
 Homophily reflects the tendency for nodes with similar characteristics or attributes to form connections and associate with each other.
@@
 To measure the presence of homophily, we measure the attribute assortativity coefficient of each of the three features (hobby, color, location).
-For a property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$, its assortativity coefficient $R$ is defined as
-$R=\frac{\sum_{i=1}^{n}M_{ii}-\sum_{i=1}^{n}a_{i}b_{i}}{1-\sum_{i=1}^{n}a_{i}b_{i}},$
-where $M$ is the mixing matrix with entries $M_{ij}$ corresponding to the fraction of edges where one end has value $P_{i}$, and the other end has value $P_{j}$, $a_{i}=\sum_{j=1}^{n}M_{ij}$ and $b_{i}=\sum_{j=1}^{n}M_{ji}$.
-The assortativity has values between $-1$ and $+1$, where a positive value denotes the tendency of the nodes to connect to other similar nodes (hence yielding a homophilous network), and a negative value corresponds to connecting to the tendency to connect to dissimilar nodes (which indicates heterophily).
-In Figure 3, we observe a consistent pattern of positive assortativity across all temperature ranges, which is higher for lower temperatures.
-Moreover, the observed assortativity at all temperatures significantly deviates from Random, which has assortativity close to 0 ($P<0.001$, t-test comparing with Random).
-Taken together, we conclude homophily emerges for the networks generated by LLMs.
-Moreover, we present the probability of each reason being mentioned in Appendix C.
+For each property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$ (indexed by $k$ or $l$), its assortativity coefficient $R$ is defined as
+$R=\frac{\sum_{k=1}^{K}M_{kk}-\sum_{k=1}^{K}a_{k}b_{k}}{1-\sum_{k=1}^{K}a_{k}b_{k}}.$
+Here $M$ represents the mixing matrix.
+Its elements $M_{kl}$ reflect the proportion of edges connecting two nodes with values $P_{k}$ and $P_{l}$, respectively.
+We define $a_{k}=\sum_{l=1}^{K}M_{kl}$ and $b_{k}=\sum_{l=1}^{K}M_{lk}$.
+Assortativity ranges from $-1$ to $+1$.
+A positive assortativity indicates nodes preferentially connect to similar ones, forming a homophilous network.
+Conversely, a negative assortativity suggests connections primarily occur between dissimilar nodes, indicating heterophily.
+In Figure 3(a), we observe a consistent pattern of positive assortativity across all temperature ranges, which is higher for lower temperatures.
+Moreover, the observed assortativity at all temperatures significantly differs from Random, which has assortativity close to 0 ($P<0.001$, t-test).
+Moreover, we present the probability of each reason being mentioned in SI Section 3.
 We find that shared hobbies appear to be much more important than shared location or favorite color.
 This is consistent with our intuition that, similarly to humans, LLMs care more about hobbies than other superficial attributes when forming social connections.
-2.2. Macro-Level Principles
+In token, we conclude homophily emerges for the networks generated by LLMs.
+Macro-Level Principles
+(a) Assortativity Coefficients
+(b) Network Instances
+(c) Communities Statistics
+Figure 3.
+Results for Principle 3 (homophily).
+(a) Figure 3(a) shows the assortativity coefficients for each network attribute created based on Principle 3.
+Each node can form at most $\delta=5$ links with the other nodes based on their attributes.
+We observe that for all temperatures, we have positive assortativity ($P<0.001$, t-test with 0), which indicates homophily.
+The assortativities are significantly different for all temperatures from random or heterophilous ($P<0.001$, t-test comparing with Random).
+(b) Networks created based on Principle 3.
+Each node can form at most $\delta=5$ links with the other nodes based on their attributes.
+The edge widths correspond to the number of common attributes shared between the nodes.
+The node colors correspond to the communities found by running the Louvain algorithm.
+(c) Louvain modularity, number of communities, and community sizes of networks created based on Principle 3.
+Lower temperatures yield higher modularity ($P<0.001$, t-test) and fewer and larger communities ($P<0.001$, t-test).
+For smaller temperatures, the communities are more “well-separated”.
+The error bars correspond to 95% confidence intervals.
 Principle 4: Community Structure
@@
 Identifying community structures is crucial for understanding the overall dynamics of a network, as it reveals patterns of relationships and interactions that might not be apparent at the global level.
-Both triadic closure and homophily contribute to the formation of community structures.By examining how these two factors contribute to network formation, we aim to gain insights into the underlying mechanisms driving community dynamics in LLM-generated networks.
-We employ the simulation results presented in Section 2.1 to determine whether community structure in networks generated by LLMs emerges from triadic closure or homophily.
+Both triadic closure and homophily contribute to the formation of community structures.
+By examining how these two factors contribute to network formation, we aim to gain insights into the underlying mechanisms driving community dynamics in LLM-generated networks.
+We employ the simulation results presented in the synthetic networks to determine whether community structure in networks generated by LLMs emerges from triadic closure or homophily.
 First, we consider the networks generated in Figure 2.
@@
 This indicates that most edges are within the same community, strengthening the community structure.
-Next, we use modularity maximization (Blondel et al., 2008) to examine the community structure that arises due to homophily.
-Intuitively, the modularity measure compares the number of edges inside a community with the expected number of edges that one would find in the cluster if the network were a random network with the same number of nodes and where each node keeps its weighted degree, but edges are otherwise randomly attached according to the Chung-Lu model (Chung and Lu, 2004).
-This random null model implicitly assumes that each node can attach to another network node.
-Formally, for a graph with weights $w_{ij}$ on its edges, the weighted modularity $Q$ (Clauset et al., 2004) of a partition with $C$ communities given by
-$Q=\sum_{c=1}^{C}\left[\frac{L_{c}}{W}-r\left(\frac{k_{c}}{2W}\right)^{2}\right],$
-Here $W$ is the sum of all weights, $L_{c}$ is the total weight of intra-community links for community $c$, $k_{c}$ is the total weighted degree of the nodes in the community $c$ and $r$ is the resolution parameter which (we take $r=1$ from now on).
-A large positive value (such as greater than 0.5) of the modularity corresponds to evidence of community structure, corresponding to deviation from the random null model.
-Maximizing $Q$ is NP-Hard, we use the Louvain algorithm (Blondel et al., 2008) to get a lower bound on the maximum modularity.
-In Figure 3(b), we illustrate the communities identified by the Louvain algorithm at various temperatures.
-Alongside, we present the average Louvain modularity values over ten simulations, with 95% confidence intervals (refer to Figure 3(c)).
-For the network’s weights, we use the number of common attributes shared between each pair of nodes.
-This is defined as $w_{ij}=\left|\left\{k:x_{i}^{(k)}=x_{j}^{(k)}\right\}\right|$ for each link $(i,j)$ in the final network.
+Next, we investigate the community structure resulting from homophily using modularity maximization (Blondel et al., 2008) (Figure 3).
+Modularity quantifies the discrepancy between the actual number of edges within communities and the expected number in a random network with identical node count and degree distribution, following the Chung-Lu model (Chung and Lu, 2004).
+This model presumes that nodes maintain their weighted degree, with edges randomly distributed.
+The weighted modularity $Q$ (Clauset et al., 2004) for a graph with edge weights $w_{ij}$ and $C$ communities is defined as
+$Q=\sum_{c=1}^{C}\left[\frac{L_{c}}{W}-r\left(\frac{k_{c}}{2W}\right)^{2}\right].$
+Here $W$ represents the total edge weights, $L_{c}$ the intra-community link weights for community $c$, $k_{c}$ the total weighted degree within community $c$, and $r$ the resolution parameter, set to 1 for our analysis.
+High modularity values (e.g., greater than 0.5) indicate significant community structuring, diverging from the random model.
+[^8] Given the NP-Hard nature of maximizing $Q$, we employ the Louvain algorithm (Blondel et al., 2008) to approximate the highest possible modularity.
+For the network’s weights, we use the number of common attributes shared between each pair of nodes: $w_{ij}=\left|\left\{k:x_{i}^{(k)}=x_{j}^{(k)}\right\}\right|$ for each link $(i,j)$ in the final network.
 Here, $x_{i}^{(k)}$ and $x_{j}^{(k)}$ correspond to the $k$-th features of $x_{i}$ and $x_{j}$, respectively.
-We observe that the Louvain partitions demonstrate community structure with positive modularity across all temperatures.
-This is confirmed by a t-test comparing the modularity to $Q=0$ of a random graph ($P<0.001$).
-The modularity is most pronounced at lower temperatures.
-As the temperature increases, the Louvain modularity also increases ($P<0.001$, t-test).
-Furthermore, lower temperatures result in fewer (and larger) communities on average, as identified by the Louvain algorithm ($P<0.001$, t-test).
-Additionally, communities tend to be visually more distinct at lower temperatures.
-We propose that higher temperatures reflect increased randomness in the decisions.
-Our results demonstrate that community structures manifest in networks generated by LLMs.
-Specifically, these networks adhere to the principle of triadic closure, which further reinforces the formation of community structures.
-Additionally, LLM-generated networks exhibit homophily, facilitating the emergence of distinct communities.
-In Section 2.3, we further illustrate that community structures are reinforced in real-world networks, where LLMs consider a combination of factors, including homophily and triadic closure, in their decisions to form connections.
+In Figure 3(b), various colors represent the communities identified by the Louvain algorithm at different temperatures.
+Notably, communities appear more distinct at lower temperatures, likely due to reduced randomness in decision-making at these temperatures.
+Additionally, Figure 3(c) presents the distribution of Louvain modularity values across simulations, indicating consistent community structure with positive modularity at all temperatures, confirmed by a t-test against a modularity of $Q=0$ for a random graph ($P<0.001$).
+Moreover, lower temperatures lead to fewer, larger communities as per the Louvain algorithm’s findings ($P<0.001$, t-test).
+Our results demonstrate that community structures manifest in networks generated by LLMs, driven by both triadic closure and homophily.
 Principle 5: Small-World
@@
 (c) Network Instances for $\beta=0.75$
-(d) Estimated $\hat{\beta}$ for $\beta=0.25$
-(e) Estimated $\hat{\beta}$ for $\beta=0.5$
-(f) Estimated $\hat{\beta}$ for $\beta=0.75$
 Figure 4.
-Figures 4(a), 4(b) and 4(c): and distribution of reasoning created based on Principle 5 using the altered Watts-Strogatz Model for $n=50,k=5,\beta\in\{0.25,0.5,0.75\}$, together with plots of the average clustering coefficient $C$ and the average shortest path length $L$.
+Simulation results for Principle 5 (small world).
+Network instances for the networks created according to Principle 5 using the altered Watts-Strogatz Model for node count $n=50$, average degree $k=5$, rewriting probability $\beta\in\{0.25,0.5,0.75\}$, together with plots of the average clustering coefficient $C$ and the average shortest path length $L$.
 The comparison is made with respect to a Watts-Strogatz graph with $n=50,k=5,\beta\in\{0.25,0.5,0.75\}$.
-In most cases, we observe strong evidence that $L\sim\log(n)$ and $C\sim 1/\log(n)$.
-Figures 4(d), 4(e) and 4(f): Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with rewiring probability $\hat{\beta}$.
+The error bars correspond to 95% confidence intervals.
+(a) $\beta=0.25$
+(b) $\beta=0.5$
+(c) $\beta=0.75$
+(d) $\beta=0.25$
+(e) $\beta=0.5$
+(f) $\beta=0.75$
+Figure 5.
+Fitted results for Principle 5 (small world).
+Top: Regression plots relating average shortest path length ($L$) and average clustering coefficient ($C$) with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$.
+Value $a$ in legends represents the slope of the regression lines.
+Bottom: Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with the estimated rewiring probability $\hat{\beta}$.
 We report the $P$-values of the t-test comparing the average shortest path length of the LLM-generated networks and the average shortest path length of the Watts-Strogatz graphs with rewiring probability $\hat{\beta}$.
+*: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.
 The small-world phenomenon is characterized by networks where nodes are interconnected in tight clusters, yet the average distance between any two nodes remains relatively short, typically scaling logarithmically with the network size (Watts and Strogatz, 1998; Kleinberg, 2000).
 This balance between high clustering and short path lengths characterizes small-world networks.
-Formally, a small-world network is a network where the average shortest path length $L$ grows proportionally to the network size $n$, [^8] This follows the definition in (Jacobs and Watts, 2021).
-namely
+A small-world network is defined by its average shortest path length $L$, which grows logarithmically with the size of the network $n$, [^9] As per the definition in (Jacobs and Watts, 2021).
+expressed as
 $L\sim\log(n).$
@@
 We employ a modified version of the model, where edge rewiring is informed by LLM queries, based on the current network structure.
-The generation process is parametrized by the number of nodes ($n$), the number of long-range connections ($k$), and the rewiring probability ($\beta$).
-The algorithm proceeds as follows:
-  1. (1)
-    Similarly to Watts-Strogatz, we first create a ring network with $n$ nodes.
-After that, for each node $[n]$, we create $k$ edges where $k/2$ edges connect to its rightmost neighbors and $k/2$ edges connect to its leftmost neighbors.
-  2. (2)
-    To create $G_{t}$, for each node $[n]$, we take its $k/2$ rightmost neighbors and rewire them with probability $\beta$.
-For each of the $k/2$ rightmost neighbors that are to be rewired, we make one query to the LLM, which indicates how the edge will be rewired.
-The choice is made by providing the LLM with all the network nodes and each node’s neighbors (i.e., the network structure).
-The model closely resembles the Watts-Strogatz model, with the primary distinction being the method of edge rewiring.
-Instead of randomly selecting edges for rewiring, as in the Watts-Strogatz model, we determine the rewiring of an edge by inquiring about the LLM and providing it with the current network structure.
-We generate networks with varying sizes ($n=10$ to $n=100$) and examine the relationships between the average shortest path length ($L$) and the average clustering coefficient ($C$) with respect to the network size ($n$).
-We test $\beta\in\{0.25,0.5,0.75\}$ and set $k=5$ as an illustration.
-In addition, we compare with the results reported in (Jacobs and Watts, 2021).
-Some instances of these networks are shown in Figures 4(a), 4(b) and 4(c).
-We run regressions relating to the average shortest path length with $\log(n)$ and observe that for all temperatures, we obtain statistically significant results (see the significance indicators in Figure 5, where most regresssions have $P<0.001$) being that the average shortest path length scales proportionally to $\log(n)$.
-Similarly, we show that the average clustering coefficient scales as $1/\log(n)$ (again, most regression results have $P<0.001$).
-Both results follow the guideline from (Jacobs and Watts, 2021) which tested the small-world properties of organizational networks.
-(a) $\beta=0.25$
-(b) $\beta=0.5$
-(c) $\beta=0.75$
-Figure 5.
-Regression plots relating average shortest path length and average clustering coefficient with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$.
-$a$ represents the slope of the regression lines.
-*** indicates $P<0.05$, ***** * indicates $P<0.01$, and ******* * * indicates $P<0.001$.
-However, a direct comparison with the Watts-Strogatz model reveals that the LLM-generated networks do not exactly mirror the statistics of Watts-Strogatz networks with corresponding $\beta$ values.
-As Figures 4(a), 4(b) and 4(c) show, we only observe weak evidence that the final generated networks have a similar average shortest path length as Watts-Strogatz with $\beta\in\{0.25,0.5,0.75\}$ ($P<0.1$, t-test comparing the resulting average shortest path lengths of the networks generated by multi-LLMs and the Watts-Strogatz average shortest path lengths for $\beta\in\{0.25,0.5,0.75\}$) and a larger average clustering coefficient than Watts-Strogatz ($P<0.1$, t-test comparing the resulting average clustering coefficient of the networks generated by multi-LLMs and the Watts-Strogatz average clustering coefficient for $\beta\in\{0.25,0.5,0.75\}$).
-To quantify this resemblance, we conduct a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks.
-The results, depicted in Figures 4(d), 4(e) and 4(f), show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings.
-In Figures 4(d), 4(e) and 4(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$, each simulation and each temperature together with the $P$-values resulting from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$.
+The generation process is parametrized by the number of nodes ($n$), average degree ($k$), and the rewiring probability ($\beta$).
+See details in Methods and Materials.
+We generated networks of various sizes, ranging from $n=10$ to $n=100$, to explore the relationship between the network size ($n$) and two key metrics: the average shortest path length ($L$) and the average clustering coefficient ($C$).
+For this analysis, we considered values of $\beta$ set at 0.25, 0.5, and 0.75, with a fixed $k=5$ to serve as a consistent parameter.
+Visual representations and average clustering coefficients are presented in Figures 4(a), 4(b) and 4(c).
+However, when directly compared with the Watts-Strogatz model, the networks generated by the LLMs do not precisely replicate the characteristics of Watts-Strogatz networks for the corresponding the rewiring probabilities ($\beta$).
+As illustrated in right panels of Figures 4(a), 4(b) and 4(c), we find weak evidence that the LLM-generated networks share the same average shortest path length as the Watts-Strogatz model for the rewiring probabilities ($\beta$) of 0.25, 0.5, and 0.75, with $P<0.1$ from a t-test comparing the average shortest path lengths.
+Additionally, LLM-generated networks exhibit a larger average clustering coefficient than those of the Watts-Strogatz model for the same rewiring probabilities $\beta$, also with a significance level of $P<0.1$ in the t-test comparisons.
+This discrepancy suggests the differences in the network structure and connectivity patterns between the LLM-generated networks and the classical Watts-Strogatz model.
+We also provide regressions analysis by examining the correlation between the average shortest path length and average clustering coefficient versus $\log(n)$ (refer to Figure 5(a), Figure 5(b), and Figure 5(c)).
+We found that across all tested temperatures, the relationships were statistically significant, with most regressions yielding $P<0.001$.
+This indicates that the average shortest path length increases proportionally with $\log(n)$.
+Similarly, for the average clustering coefficient, we demonstrated that it inversely scales with $1/\log(n)$, with the majority of regression analyses also showing $P<0.001$.
+These findings align with the small-world properties of organizational networks as documented in the study by (Jacobs and Watts, 2021), suggesting that these characteristics are not only prevalent but also predictable across different network sizes.
+To quantify how LLM-generated networks resemble Watts-Strogatz networks, we fit the estimated $\hat{\beta}$ values for each LLM-generated network.
+[^10] We conducted a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks.
+In Figures 5(d), 5(e) and 5(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$ and each temperature.
+Here $P$-values result from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$.
+These results show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings.
 In conclusion, our analysis demonstrates that LLM-generated networks exhibit key small-world properties, with logarithmic scaling of average shortest path lengths and inverse logarithmic scaling of average clustering coefficients.
-While these networks do not perfectly align with the Watts-Strogatz model, they exhibit similar structural characteristics, underscoring the potential of LLMs in generating small-world networks.
-2.3. Real-World Networks
+While these networks do not perfectly align with the Watts-Strogatz model, they exhibit similar structural characteristics.
+Decisions on Real-World Networks
 To investigate the behavior of LLMs in real-world network formation contexts, we employ the Facebook100 dataset (Traud et al., 2012) and a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972).
 This dataset comprises “friendship” networks from one hundred American colleges and universities, captured at a specific moment in time from Facebook’s online social network.
-We model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives.
-2.3.1. Discrete Choice Model
-For each node $i_{t}$ that we consider at time $t$, we randomly remove one of its current friends from the real-world network.
-After we remove a neighbor for each of $i_{1},\dots,i_{T}$, we end up with the network $G_{1}$, which we use as a seed network for the LLM agents.
-Subsequently, during the link formation process, we present each node $i_{t}$ with a set of candidate nodes (denoted by $A_{t}$), comprising one of the previously removed friends and other nodes that are not their friends.
-We then instruct the LLM to form a link with one of the candidates, providing the attributes of the candidates and the social network structure to aid its decision-making.
-These choices are made sequentially.
-We use the utility of the model for each node for each sequential decision of network formation:
-$U_{ij,t}=\theta_{\mathrm{PA}}\log d_{j,t}+\theta_{\mathrm{H}}\log w_{ij}+\theta_{\mathrm{TC}}\log c_{ij,t}+\epsilon_{ij,t}.$
-In this equation, $\theta_{\mathrm{PA}}$ measures the strength of preferential attachment based on the degree $d_{j,t}$ of $j$ at step $t$, $\theta_{\mathrm{H}}$ measures the strength of homophily based on the similarity $w_{ij}$ (i.e.
-number of common attributes) between $i$ and $j$, and $\theta_{\mathrm{TC}}$ measures the strength of triadic closure, based on the number of common neighbors $c_{ij,t}$ between $i$ and $j$ at step $t$.
-The error term $\epsilon_{ij,t}$ is distributed as i.i.d.
-standard Gumbel [^9] The standard Gumbel distribution has CDF $e^{e^{-x}}$..
-All variables are first normalized based on their range, and then the log transformation is taken.
-The multinomial logit model (MNL) indicates that the probability that $i$ links to $j$ at step $t$ is given by
-$p_{ij,t}=\Pr\left[\mathrm{argmax}_{r\in A_{t}}U_{ir,t}=j\right]=\frac{d_{j,t}^{\theta_{\mathrm{PA}}}w_{ij}^{\theta_{\mathrm{H}}}c_{ij,t}^{\theta_{\mathrm{TC}}}}{\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{ir}^{\theta_{\mathrm{H}}}c_{ir,t}^{\theta_{\mathrm{TC}}}}.$
-Given a sequence of nodes $i_{1},\dots,i_{T}\in V$ and choices (denoted by subscripted $j$) $j_{1}\in A_{1},\dots,j_{T}\in A_{T}$, the parameters can be found by maximizing the log-likelihood function.
-To get the standard errors of the coefficients and the corresponding $P$-values, we follow the process outlined in (Overgoor et al., 2019).
-2.3.2. Experiment Setup and Rationale
-We use datasets that refer to three US colleges – Caltech36, Swarthmore42, and UChicago30. They represent different sizes of networks: $n=769$, $n=1,659$, and $n=6,591$ nodes, respectively.
-For the Caltech36 ($n=769$) and Swarthmore42 ($n=1,659$) datasets, we consider all of the nodes of the networks as nodes that have to make a choice.
+We model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives (see Methods and Materials).
+We use datasets that refer to three US colleges – Caltech36, Swarthmore42, and UChicago30.
+They represent different sizes of networks: $n=769$, $n=1,659$, and $n=6,591$ nodes, respectively.
+For the Caltech36 ($n=769$) and Swarthmore42 ($n=1,659$) datasets, we consider all of the network nodes to be nodes that have to make a choice.
 For the UChicago30 dataset, we consider a randomly sampled subset of 2,000 nodes because of the limited context window of GPT models.
@@
 Figure 6.
+Regression results for real-world network experiments.
 Multinomial logit coefficients for three networks from the Facebook100 dataset.
 The standard errors of the estimates are shown with error bars.
-The full regression tables can be found in Appendix A.
+The full regression tables can be found in SI Section 1.
 All coefficients are statistically significant, where the null hypothesis corresponds to the respective parameter being equal to 0.
-2.3.3. Results
-We present the regression results in Figure 6101010More detailed results can be found in Table S.1 in Appendix A.
+We regress network formation decisions on standardized scores reflecting the three micro-level principles.
+We present the regression results in Figure 6.111111More detailed results can be found in Table S1 in SI Section 1.
 Our analysis reveals that the degree, the number of common neighbors, and the number of shared interests are all statistically significant factors influencing the choices of LLMs ($P<0.001$) across all datasets.
@@
 Specifically, the relationship $\hat{\theta}_{\mathrm{H}}>\hat{\theta}_{\mathrm{TC}}>\hat{\theta}_{\mathrm{PA}}$ holds across all examined datasets and temperature settings.
-In addition to the coefficient $\hat{\theta}_{\mathrm{TC}}$ of triadic closure, we measure the percentage change in the average shortest path length and the average clustering coefficient between the network before the edge deletions and the network after the choices made by the LLM agents (see Table S.1).
-We observe that the average clustering coefficient also decreases compared to the initial network, indicating the weakening of triadic closure compared to the initial network.
-We also provide reasoning analysis.
-Figure S.8 demonstrates that the rationales provided by LLMs aligns with the aforementioned findings.
-In particular, LLMs emphasize homophily, indicating that nodes sharing similar characteristics—such as student/faculty status, high school, and graduation year—are more likely to form connections.
-Conversely, the effects of triadic closure and preferential attachment are less prominent.
-We also compare LLMs’ prediction with random guessing under a link prediction framework.
-We evaluate the percentage increase in accuracy, where accuracy is defined as the likelihood of selecting a previously existing neighbor from a set of alternatives, compared to making a random selection from 20 possible options.
-As shown in Appendix A, LLMs demonstrate a substantially higher accuracy in their selections than would be expected from random guessing.
-Moreover, we examine whether the community structure is reinforced after the network formation decision made by LLMs.
-We run the Louvain algorithm 10 times for both the initial graph $G_{1}$ and the final graph $G_{T}$ and perform a t-test between the resulting modularities (see Appendix A).
+Further insights are provided in SI Section 4, where we provide the reasoning analysis.
+The analysis points to a marked emphasis on homophily, indicating that nodes with similar characteristics—such as student/faculty status, high school, and graduation year—are more likely to connect.
+In contrast, triadic closure and preferential attachment have a relatively lesser impact.
+Our study also includes a comparison of LLMs’ predictive capabilities against random guessing in a link prediction framework.
+We assess the percentage increase in accuracy, defining accuracy as the probability of selecting an existing neighbor from a set of alternatives, as opposed to a random selection from 20 possible options.
+As detailed in SI Section 1, LLMs significantly outperform random guessing in accuracy.
+Furthermore, we investigate whether LLMs reinforce community structure post-network formation.
+We run the Louvain algorithm 10 times for both the initial graph $G_{1}$ and the final graph $G_{T}$ and perform a t-test between the resulting modularity scores (see SI Section 1).
 We observe that the modularity increases in all cases and is statistically significant.
 This suggests that LLMs make choices that naturally yield a reinforcement of the community structure.
-To reflect the relation with “small-world” phenomenon, we measure the percent change in the average shortest path length between the network before the edge deletions and the network after the choices made by the LLM agents (see Appendix A).
-We observe a small decrease in the average path length in the new network in almost all cases indicating that the average shortest path length decreases slightly, indicating strengthening of the small-world phenomenon.
-Taken together, our analysis indicates that although preferential attachment, homophily, and triadic closure are all present in the network formation behaviors of LLMs, homophily emerges as the predominant driving force.
-Moreover, both community structure and small-world phenomenon are reinforced.
-3. Discussion
+Additionally, to explore the “small-world” phenomenon, we measure the percent change in average shortest path length between the pre- and post-decision networks (detailed in SI Section 1).
+Our findings reveal a minor decrease in average path length in most cases, suggesting a slight strengthening of the small-world phenomenon.
+In summary, our analysis demonstrates that while homophily, triadic closure, and preferential attachment are integral to the network formation behaviors of LLMs, homophily is the dominant factor.
+Furthermore, both community structure and the small-world phenomenon are reinforced after network formation decisions of LLMs.
+Discussion
 Our study conducted a comprehensive evaluation of LLMs’ network formation preferences, encompassing both micro-level network principles such as preferential attachment, triadic closure, and homophily and macro-level network principles like community structure and the small-world phenomenon.
@@
 Furthermore, through discrete choice modeling, we explored the emergence of these properties in real-world networks.
-Our results reveal that the agents’ selections are predominantly driven by homophily, followed by triadic closure, and then preferential attachment.
+Our real-world network results reveal that the agents’ selections are predominantly driven by homophily, followed by triadic closure, and then preferential attachment.
 Additionally, we observed a strengthening of community structure in real-world network experiments.
@@
 This insight contributes to developing socially aware LLMs that align with human preferences, with potential applications ranging from chatbots and virtual assistants to collaborative systems and social media platforms.
-Furthermore, our study may pave the way for innovative applications of LLMs in network science research.
+Furthermore, our study may pave the way for innovative applications of LLMs in network science and computational social science research.
 Given that LLMs adhere to fundamental network principles and exhibit behaviors akin to human network formation, they hold significant potential for use in simulations of social systems and agent-based modeling.
-LLM agents have the capability to produce more sophisticated and nuanced behaviors, including negotiation, collaboration, and competition.
-This opens up opportunities to simulate networked environments such as marketplaces or organizational structures, where LLMs can model decision-making processes and interaction patterns.
+As LLM agents gain increasing ability to exhibit more complex and nuanced behaviors, such as negotiation, collaboration, and competition, this presents opportunities to replicate networked environments like marketplaces or organizational structures.
+In these settings, LLMs could effectively model decision-making processes and interaction dynamics.
 Looking ahead, we propose three promising directions for future research.
@@
 1. Methods
-4.1. Experimental Procedure
+Methods and Materials
+# Task Your task is to select a set of people to be friends with.
+# Profile Your profile is given below after chevrons: <PROFILE>$F(\{i_{t}\})$</PROFILE> # Candidate Profiles The candidate profiles to be friends with are given below after chevrons: <PROFILES>$F(A_{t})$</PROFILES> # Output The output should be given a list of JSON objects with the following structure [ {{ "name" : name of the person you selected, "reason" : reason for selecting the person }}, ...
+] # Notes - The output must be a list of JSON objects ranked in the order of preference.
+- You can make at most $\delta$ selections.
+ALGORITHM 1 Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$.
+Experimental Procedure
 In our study, we performed experiments to assess whether key network principles at both the micro-level (such as preferential attachment, triadic closure, and homophily) and the macro-level (including community structure and weak ties) align with classical network models.
 Subsequently, we utilized real-world networks to determine the factors that are most heavily weighted by LLMs.
-Our experiments span a time series of $T$ steps, with a sequence of network structures denoted as $G_{1},G_{2},\dots,G_{T}$.
+Our experiments span a time series of $T$ steps, with a sequence of network structures denoted as $G_{1},G_{2},\dots,G_{T}$ with vertex sets $V_{1},\dots,V_{T}$.
 The initial network, $G_{1}$, is referred to as the seed network.
@@
 This approach is adopted because LLMs have demonstrated proficiency in processing code-like structures, such as HTML and JSON.
-4.2. Summarizing the Reasoning of the LLM Agents
+Summarizing the Reasoning of the LLM Agents
 In addition to analyzing the innate preferences of LLM agents, we employ an interview-like approach to probe the rationale behind their decisions.
 Following the LLMs’ decisions on network formation, we inquire about their reasoning for making those choices.
-We then aggregate all the provided rationales using GPT models (GPT-3.5 for Section 2.1 and GPT-4 for Section 2.3) by asking it to identify the common themes, resulting in the creation of multiple categories.
+We then aggregate all the provided rationales using GPT models (GPT-3.5 for the synthetic networks and GPT-4 for the real-world networks) by asking it to identify the common themes, resulting in the creation of multiple categories.
 To determine the category to which each LLM’s reasoning belongs, we subsequently ask the GPT models to classify each rationale into the appropriate category.
-4.3. Estimating the Parameters of the Discrete Choice Model
+Details for Small-World Experiments
+The algorithm of generating LLM networks is described as follows:
+  1. (1)
+    Similarly to Watts-Strogatz, we first create a ring network with $n$ nodes.
+After that, for each node $[n]$, we create $k$ edges where $k/2$ edges connect to its rightmost neighbors and $k/2$ edges connect to its leftmost neighbors.
+  2. (2)
+    To create $G_{t}$, for each node $[n]$, we take its $k/2$ rightmost neighbors and rewire them with probability $\beta$.
+For each of the $k/2$ rightmost neighbors that are to be rewired, we make one query to the LLM, which indicates how the edge will be rewired.
+The choice is made by providing the LLM with all the network nodes and each node’s neighbors (i.e., the network structure).
+The model closely resembles the Watts-Strogatz model, with the primary distinction being the method of edge rewiring.
+Instead of randomly selecting edges for rewiring, as in the Watts-Strogatz model, we determine the rewiring of an edge by inquiring about the LLM and providing it with the current network structure.
+The Discrete Choice Model in Real-World Network Experiments
+For each node $i_{t}$ that we consider at time $t$, we randomly remove one of its current friends from the real-world network.
+After we remove a neighbor for each of $i_{1},\dots,i_{T}$, we end up with the network $G_{1}$, which we use as a seed network for the LLM agents.
+Subsequently, during the link formation process, we present each node $i_{t}$ with a set of candidate nodes (denoted by $A_{t}$), comprising one of the previously removed friends and other nodes that are not their friends.
+We then instruct the LLM to form a link with one of the candidates, providing the attributes of the candidates and the social network structure to aid its decision-making.
+These choices are made sequentially.
+We use the utility of the model for each node for each sequential decision of network formation:
+$U_{ij,t}=\theta_{\mathrm{PA}}\log d_{j,t}+\theta_{\mathrm{H}}\log w_{ij}+\theta_{\mathrm{TC}}\log c_{ij,t}+\epsilon_{ij,t}.$
+In this equation, $\theta_{\mathrm{PA}}$ measures the strength of preferential attachment based on the degree $d_{j,t}$ of $j$ at step $t$, $\theta_{\mathrm{H}}$ measures the strength of homophily based on the similarity $w_{ij}$ (i.e.
+number of common attributes) between $i$ and $j$, and $\theta_{\mathrm{TC}}$ measures the strength of triadic closure, based on the number of common neighbors $c_{ij,t}$ between $i$ and $j$ at step $t$.
+The error term $\epsilon_{ij,t}$ is distributed as i.i.d.
+standard Gumbel.
+[^12] The standard Gumbel distribution has CDF $e^{e^{-x}}$.
+All variables are first normalized based on their range, and then the log transformation is taken.
+The multinomial logit model (MNL) indicates that the probability that $i$ links to $j$ at step $t$ is given by
+$p_{ij,t}=\Pr\left[\mathrm{argmax}_{r\in A_{t}}U_{ir,t}=j\right]=\frac{d_{j,t}^{\theta_{\mathrm{PA}}}w_{ij}^{\theta_{\mathrm{H}}}c_{ij,t}^{\theta_{\mathrm{TC}}}}{\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{ir}^{\theta_{\mathrm{H}}}c_{ir,t}^{\theta_{\mathrm{TC}}}}.$
+Given a sequence of nodes $i_{1},\dots,i_{T}\in V$ and choices (denoted by subscripted $j$) $j_{1}\in A_{1},\dots,j_{T}\in A_{T}$, the parameters can be found by maximizing the log-likelihood function.
+To get the standard errors of the coefficients and the corresponding $P$-values, we follow the process outlined in (Overgoor et al., 2019).
+Estimating the Parameters of the Discrete Choice Model
 To estimate the parameters of the discrete choice model, we optimize the following log-likelihood function
-$(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})=\operatorname*{arg\,max}_{(\theta_{\mathrm{PA}},\theta_{\mathrm{TC}},\theta_{\mathrm{H}})\in\mathbf{R}^{3}}\sum_{t=1}^{T}\left(\theta_{\mathrm{PA}}\log d_{j_{t},t}+\theta_{\mathrm{H}}\log w_{i_{t}j_{t}}+\theta_{\mathrm{TC}}\log c_{i_{t}j_{t},t}-\log\left(\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{i_{t}r}^{\theta_{\mathrm{H}}}c_{i_{t}r,t}^{\theta_{\mathrm{TC}}}\right)\right),$
+$(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})=\operatorname*{arg\,max}_{(\theta_{\mathrm{PA}},\theta_{\mathrm{TC}},\theta_{\mathrm{H}})\in\mathbf{R}^{3}}\sum_{t=1}^{T}\bigg{(}\theta_{\mathrm{PA}}\log d_{j_{t},t}+\theta_{\mathrm{H}}\log w_{i_{t}j_{t}}+\theta_{\mathrm{TC}}\log c_{i_{t}j_{t},t}-\log\big{(}\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{i_{t}r}^{\theta_{\mathrm{H}}}c_{i_{t}r,t}^{\theta_{\mathrm{TC}}}\big{)}\bigg{)},$
 where $i_{1},\dots,i_{T}$ are the chooser nodes (i.e., the LLM agents who want to form a link), and $j_{1},\dots,j_{T}$ are the nodes which are chosen from the alternative sets $A_{1},\dots,A_{T}$.
@@
 (Overgoor et al., 2019; Train, 2009)).
+Data and Code Availability
+Data and code are available at GitHub: https://github.com/papachristoumarios/llm-network-formation.
 Acknowledgements.
-M.P. is supported by a scholarship from the Onassis Foundation (Scholarship ID: F ZT 056-1/2023-2024).
+The authors would like to thank Jon Kleinberg, Yanbang Wang, Yuanqi Du, the attendees of the Learning on Graphs NYC meetup, and the participants of Cornell AI, Policy, and Practice working group for their valuable feedback on the current version of the paper.
+M.P.
+is supported by a scholarship from the Onassis Foundation (Scholarship ID: F ZT 056-1/2023-2024).
 Appendix A Regression Table
 In Table S.1, we report the regression coefficient for the regression in the real-world network data.
-The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 2.3 (also shown in Figure 6) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0).
+The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 1.C (also shown in Figure 5) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0).
 Next, LL corresponds to the log-likelihood of the fitted model, and AIC corresponds to the Akaike Information Criterion.
@@
 Acc. = Accuracy, $L$ = Average Path Length, $C$ = Average Clustering Coefficient, $\Delta Q$ (t-stat) = Modularity change t-statistic
-:*${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
+${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
 Table S.1.
@@
 1.5 0.36*** (0.16) 1.72*** (0.31) 0.55*** (0.20) -1,522.47 3,052.94
-Note :*${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
+Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
 Table S.2. Multinomial logit coefficients for Caltech36. The standard errors of the estimates are shown in parentheses.
@@
 1.5 0.19*** (0.14) 1.50*** (0.33) 0.59*** (0.10) -3,139.42 6,286.83
-Note :*${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
+Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
 Table S.3. Multinomial logit coefficients for Swarthmore42. The standard errors of the estimates are shown in parentheses.
@@
 1.5 0.22*** (0.12) 1.78*** (0.27) 0.41*** (0.07) -2,033.49 4,074.98
-Note :*${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
+Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
 Table S.4. Multinomial logit coefficients for UChicago30. The standard errors of the estimates are shown in parentheses.
@@
 After LLMs form networks, we ask them about their reasoning.
-We then use GPT models (GPT-3.5 for Section 2.1 and GPT-4 for Section 2.3) to group their explanations into common themes, creating various categories.
+We then use GPT models (GPT-3.5 for Section 1.A.
+and GPT-4 for Section 1.C) to group their explanations into common themes, creating various categories.
 To classify each LLM’s reasoning, we have the GPT models assign them to the relevant category.
 C.1. Micro-level Principles
-In Figure S.7, we show the reasoning of the LLM agents from the networks of Section 2.1.
-In Figures 7(a) and 7(b), the reasons mean the following:
-  • Well-Connectedness: it corresponds to reasons related to the person being well-connected and having many friends.
-  • Mutual Friends: it is related to having many friends in common with the person so that the person can introduce them to their friends.
+In Figure S.7, we show the reasoning of the LLM agents from the networks of Section 1.A.
 We observe that for Figure 7(a), the well-connectedness reason becomes less important as the temperature increases while having more mutual friends becomes more important as the temperature increases.
@@
 A possible explanation for this asymmetry is the fact that different types of features are provided, i.e., in the case of Figure 7(a) the neighbors of each node are provided, whereas in the case of Figure 7(b) the common neighbors with the query node $i_{t}$ are provided.
-Finally, Figure 7(c) contains the reasons for Figure 3(a) based on having a common hobby, a common color, or a common location.
+Finally, Figure 7(c) contains the reasons for Figure 3 based on having a common hobby, a common color, or a common location.
 We observe that the LLM agents make choices primarily based on sharing the same hobby and then based on a common location or a common color (on an almost equal basis).
@@
 Figure S.7.
-Distribution of reasoning across categories for the networks of Figures 1(a), 2(a) and 3(b).
+Distribution of reasoning across categories for the networks of Figures 1(a), 2(a), and 3.
 The classification uses GPT-3.5 on 20 randomly sampled reasons repeated over five times (micro-averaged frequencies).

### v4

@@
 Abstract.
-Social networks shape opinions, behaviors, and information dissemination in human societies.
-As large language models (LLMs) increasingly integrate into social and professional environments, understanding their behavior within the context of social interactions and networks becomes essential.
-Our study analyzes LLMs’ network formation behavior to examine whether the dynamics of multiple LLMs are similar to or different from human social dynamics.
-We observe that LLMs exhibit key social network principles, including preferential attachment, triadic closure, homophily, community structure, and the small-world phenomenon, when asked about their preferences in network formation.
-We also investigate LLMs’ decision-making based on real-world networks, revealing that triadic closure and homophily have a stronger influence than preferential attachment and that LLMs perform well in network formation predictions.
-Overall, our study opens up new possibilities for using LLMs in network science research and helps develop socially aware LLMs by shedding light on their social interaction behaviors and exploring their impacts on social dynamics.
+Social networks fundamentally shape human opinions, behaviors, and the dissemination of information.
+As large language models (LLMs) like GPT, Claude, and Llama increasingly integrate into social and professional settings, understanding their behavior in the context of social interactions and network formation becomes essential.
+This study develops a framework to systematically examine whether the network formation behaviors of multiple LLMs approximate certain aspects of human network dynamics.
+By simulating interactions among LLM agents across various model families, we observe that these models consistently exhibit key patterns associated with social network principles—including preferential attachment, triadic closure, homophily, community structure, and the small-world phenomenon—when forming networks.
+Moreover, LLMs adapt their network formation strategies based on each network’s characteristics, reflecting the context-dependent nature of human behavior: in Facebook networks, they prioritize triadic closure and homophily, mirroring close-knit friendships; in phone networks, homophily and preferential attachment dominate, capturing personal and professional connections, while in employment networks, LLMs favor heterophily and high-degree connections, aligning with career advancement dynamics.
+These results open new avenues for using LLMs in network science research, with potential applications in agent-based modeling and synthetic network generation.
 large language models, networks, network formation, link prediction, multi-agent systems
@@
 Introduction
-Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023) and Llama 2 (Touvron et al., 2023), have shown promising developments in AI techniques and their integration into real-life applications.
+Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Llama (Touvron et al., 2023), has shown promising developments in AI techniques and their integration into real-life applications.
 It is thus crucial to comprehend AI actions to ensure they align with human expectations, mitigate potential risks, and maximize their benefits.
 Misaligned AI actions may lead to unintended consequences, such as biased decision-making, fairness issues, and the miscoordinative or non-cooperative behavior (Rahwan et al., 2019).
-Recently, researchers have started to apply social science methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods, to study LLMs.
+Recently, researchers have started to apply social science methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods (Chew et al., 2023), to study LLMs.
 These methods not only reveal the capabilities and interpretability of LLMs but also suggest their potential for applications in social science (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
@@
 LLMs have shown great potential in social contexts, notably as intelligent personal assistants that facilitate social and prosocial interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)).
-However, less is known about how LLMs’ innate behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023).
+However, less is known about how LLMs’ behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023).
 This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the networks of human relationships, which is a fundamental aspect of social systems.
@@
 Specifically, we examine micro-level social network properties including preferential attachment (Barabási and Albert, 1999), triadic closure (Granovetter, 1973), and homophily (McPherson et al., 2001), as well as macro-level properties including community structure (Newman and Girvan, 2004), and the small-world phenomenon (Kleinberg, 2000; Watts and Strogatz, 1998).
-In simulations conducted using synthetic networks, we have the following findings: First, we observe that LLMs can generate networks exhibiting preferential attachment and following scale-free degree distributions.
-Moreover, LLMs demonstrate strong tendencies toward triadic closure and exhibit tendency of homophily.
-These tendencies further explain the emergence of community structures in networks formed by multi-LLMs.
-Additionally, we show that LLM agents exhibit dynamics consistent with small-world phenomena akin to those observed in the Watts-Strogatz model (Watts and Strogatz, 1998).
-Next, we compare between these principles to identify the key factors driving network dynamics in real-world network formation contexts.
-To achieve this, we consider real-world networks and ask each LLM to simulate a node in the network.
-We provide information on their network structure and ask for their decisions for the next link formation or dissolution step.
-We approach this problem as a discrete choice modeling problem (Liben-Nowell and Kleinberg, 2003; Overgoor et al., 2019).
-In this real-world setting, we find that LLMs’ choices are primarily weighted on homophily and triadic closure, compared to preferential attachment, and that LLMs substantially exceed random guessing when predicting new links.
-Additionally, the resulting choices of LLM agents strengthen the network’s community structure.
-Our study contributes to the literature on LLMs in multiple ways.
-First, it advances the discussion on aligning LLMs with human preferences, which has potential applications in developing socially aware LLMs for real-life interactions.
-Insights into LLMs’ innate behaviors in network formation can inform the design of LLMs for effective collaboration in social and professional settings.
-Second, our study could contribute to understanding the impact of LLMs on social dynamics and norms, such as their role in shaping online communities’ interaction patterns or influencing the adoption of new communication norms in professional settings.
-Finally, our work has direct applications to synthetic dataset generation, a problem important for the network science community.
-By controlling for a few parameters, such as the temperature, our work can be utilized to generate a variety of datasets, as well as generate synthetic datasets from real-world data to protect privacy.
-(a) Structure-based
-(b) Degree-based
-Figure 1.
-Results for Principle 1 (preferential attachment): We display simulated networks comprising 200 nodes across different temperatures.
-For structure-based simulations (Figure 1(a)), the multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$, while for degree-based simulations (Figure 1(b)), node degree data $\{d_{j,t}:j\in V_{t}\}$ was provided ($V_{t}$ corresponds to the vertex set of the network $G_{t}$ at round $t$).
-The degree distributions were then compared with those of a BA network with $n=200$ nodes.
-(a): Power-law degree distributions are evident ($P>0.5$, K-S test), with the networks at a temperature of 1.5 closely resembling the BA model ($P>0.1$, K-S test).
-(b): With degree information only, the networks form more unrealistic star-like structures, diverging from scale-free configurations and more closely mirroring a core-periphery network structure.
+In synthetic network simulations, LLMs displayed preferential attachment, homophily, and triadic closure, resulting in the formation of community structures and small-world dynamics.
+More notably, in real-world social network simulations, we find that LLMs prioritized triadic closure and homophily over preferential attachment when forming new links, indicating a strong preference for connecting with similar nodes or shared acquaintances.
+Additionally, in a telecommunication network, LLMs tended to prioritize homophily and preferential attachment over triadic closure, and in a company network, the agents who corresponded to employees formed links frequently with managers, which showcases behavior that is consistent with human social mobility principles.
+Generally, LLMs not only exhibit fundamental social network formation principles in synthetic simulations but also adapt their strategies based on the context of real-world networks, mirroring human social behaviors specific to each setting.
+As LLM technology continues to evolve, our study serves as an early exploration of their potential in social network studies, with several significant implications for future studies.
+First, our study demonstrates the potential of LLMs for agent-based modeling.
+By simulating decision-making processes that approximate human-like behavior across various network settings, LLMs can provide valuable insights into the emergence of social phenomena.
+Although these models are still in early stages of development, they offer an intriguing framework for studying and designing systems that can mimic key aspects of real-world dynamics.
+This opens up possibilities for applying LLMs to explore and understand complex behaviors in social, professional, and collaborative environments.
+Second, our work highlights the potential of LLMs for synthetic dataset generation, a critical area in network science.
+Although the accuracy of LLM-based predictions is not yet perfect like all other link prediction models, this approach is particularly valuable in scenarios where privacy concerns limit access to real-world data.
+By simulating realistic datasets that capture important network properties, LLMs can facilitate research and experimentation without compromising sensitive information.
 Results
-In our study, we leveraged state-of-the-art LLMs through the OpenAI Application Programming Interface (API).
-Our methodology involved constructing multiple LLMs, each acting as an independent agent within separate conversational threads.
-Each LLM agent was prompted with network information about its peers and faced with decision-making scenarios.
-For synthetic network simulations, we employed GPT-3.5 agents, whereas GPT-4 agents were used for real-world network analysis.
-[^1] We utilized the gpt-3.5-turbo model for GPT-3.5 and the gpt-4-1106-preview model for GPT-4, both provided by OpenAI.
-Notably, both GPT-3.5 and GPT-4 yielded consistent results.
-We adopt a multi-LLM framework, emphasizing the multi-agent aspect of our experimental design, which shares similar spirits with recent research in the field (Leng et al., 2023; Akata et al., 2023; Brookins and DeBacker, 2023; Aher et al., 2023; De Marzo et al., 2023).
-To assess the robustness of our findings, we varied the ‘temperature’ settings in LLMs, which modulate the output variability, in a dynamic network environment.
-Additionally, we adopted an interview-like method to probe the rationale behind the LLMs’ decision-making processes, enriching our understanding of their behavior in network simulations.
-In our analysis, we begin with an examination of three fundamental micro-level network principles: preferential attachment, triadic closure, and homophily.
-Subsequently, we shift our focus to two macro-level network phenomena: community structure and the small-world phenomenon.
-Finally, we assess the resilience of these principles as LLM agents interact with real-world networks.
-This allows us to compare and contrast the network formation preferences between LLMs and humans.
+In this study, we investigated whether LLMs exhibit fundamental principles of network formation observed in human social networks.
+By simulating multiple LLM agents acting independently within separate conversational threads, we examined their behaviors in decision-making scenarios involving network connections.
+We focused on three micro-level network principles – preferential attachment, triadic closure, and homophily – and two macro-level phenomena—community structure and the small-world effect.
+To assess the robustness of our findings, we varied the temperature settings of different LLM models, including GPT-3.5-turbo, GPT-4o Mini, Llama 3 (70b-instruct), and Claude 3.5 Sonnet.
+We also experimented with different environmental prompts (e.g., friendship, collaboration, community) to test prompt sensitivity.
+Additionally, we employed an interview-like method to probe the LLMs’ decision-making rationale and conducted experiments using Chain-of-Thought (CoT) reasoning (Wei et al., 2022) (the experiments are deferred to Appendix C).
+Finally, we extended our analysis to real-world networks, including a social media friendship network, a telecommunication network, and a company collaboration network, to compare the network formation preferences between LLMs and humans.
+More information about the experimental procedure can be found in the Methods and Materials section of this paper.
 Micro-Level Properties
 Principle 1: Preferential Attachment
+(a) Probability of connecting to top-$k$ nodes for different models, temperatures, and environments
+(b) Power law fits ($\hat{\gamma}$) and standard errors for different models, temperatures, and environments
+(c) Network instances generated by GPT-3.5 agents
+Figure 1.
+Results for Principle 1 (preferential attachment) The multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$.
+Top: Probability of connecting to top-$k$-degree nodes for varying model (temperature is fixed to 1.0 and environment to baseline), temperature (model fixed to GPT-3.5 and environment to baseline) and environment (model fixed to GPT-3.5 and environment temperature to 1.5) for networks generated according to Principle 1 with $n=200$ nodes.
+Middle: Power Law exponents and standard errors for varying model, temperature, and environment.
+Bottom: Simulated networks.
+Power-law degree distributions are evident ($P>0.5$, K-S test), with the networks at a temperature of 1.5 closely resembling the Barabási-Albert model ($P>0.1$, K-S test) for GPT-3.5 agents.
 Preferential attachment is a fundamental concept in network science, illustrating how nodes in a network gain connections over time, leading to a scale-free degree distribution characterized by a few highly connected nodes (Barabási and Albert, 1999; Bianconi and Barabási, 2001).
-In our study, we investigate whether LLMs manifest the preferential attachment mechanism.
-Starting with an empty network, we introduce nodes sequentially at each time step $t$.
-A new node $i_{t}$ is added during time step $t$, with the network’s existing nodes providing context about their connections.
-We generated networks with $n=200$ nodes, a size large enough to observe meaningful patterns in the degree distribution.
-Our simulations include two distinct types: structure-based, where the network’s structure was provided, and degree-based, where only the degree information of potential connections was given.
-[^2] For details on how features were represented in the prompts, refer to SI Section 5.
-The graphical outputs from our structure-based simulation, as illustrated in Figure 1(a), reveal a distinct pattern where a small number of nodes possess a high number of connections, whereas the majority exhibit fewer links.
-This pattern suggests a scale-free degree distribution, characterized by the probability $\pi(d)$ that a node has degree $d$, mathematically represented as:
-(1) $\pi(d)\propto d^{-\gamma},\quad\text{where }\gamma>1.$
-Interestingly, we also observe that the number of highly connected “super nodes” increases with temperature.
-[…]
-Furthermore, in degree-based simulations in a similar spirit of ((De Marzo et al., 2023); see Figure 1(b)), we noted formations resembling a “star” around central nodes.
-These distributions deviate from the scale-free patterns seen in statistical models like the BA model ($P<0.001$, two-sample K-S test), aligning more closely with core-periphery structures (Borgatti and Everett, 2000).
-To interpret the LLM agents’ decision-making, we analyzed the reasoning they provided, focusing on the structure-based results (Figure 1(a)) [^4] The degree-based results’ reasoning is not included as decisions were based solely on degree information..
-Two primary reasons emerged, varying with temperature (detailed in SI Section 3):
-  • Well-Connectedness: Relating to the node’s degree (number of connections).
-  • Mutual Friends: Relating to the number of shared connections.
-Interestingly, as the temperature increased, the prevalence of the well-connectedness rationale decreased, while the mutual friends’ rationale increased.
-In conclusion, our analysis demonstrates that LLMs, when provided with network structure, can replicate the preferential attachment phenomenon.
+To test if LLM agents exhibit preferential attachment, we simulated network growth by sequentially adding nodes to an initially empty network.
+Each new node was prompted with information about existing nodes, and the person to connect with was decided.
+We generated networks with $n=200$ nodes to observe meaningful degree distributions [^1] Note that we provide the full network structure in the prompt, so models are not inherently biased toward forming links with the highest-degree nodes..
+On a micro-scale, Figure 1(a) illustrates the probability of connecting to a top-$k$ node as a function of its degree percentile ($k/n$).
+To demonstrate the tendency toward preferential attachment, we compare these probabilities to a null model assuming random connections (represented by dashed lines), where the likelihood of connecting to a top-$k$ node is simply $k/n$.
+Our findings reveal that all models prefer connecting to higher-degree nodes.
+Notably, GPT-3.5 exhibits a weaker preference, while other, arguably more capable models, show an even stronger inclination toward preferential attachment.
+Using GPT-3.5 as an example, we examine the effect of temperature – a parameter controlling the variability of model output – on this tendency.
+At lower temperatures, the model makes fewer stochastic choices and, as a result, is more likely to connect to high-degree nodes.
+We also vary the prompt to explore the influence of “environment”-contextual settings such as school, work, or community.
+The results show slight variations compared to the baseline (GPT-3.5 with temperature = 1.5), yet the tendency for preferential attachment persists across environments.
+In all cases, the observed curves lie above the null model, underscoring the presence of preferential attachment.
+Next, we investigate the degree distribution of the resulting graphs.
+As shown in Figure 1(c), the resulting networks display a pattern where a few nodes have many connections while most have few, indicative of a scale-free distribution, with form:
+(1) $\pi(d)\propto d^{-\gamma},\quad\text{where}\quad\gamma>1.$
+We estimated the exponent $\gamma$ for different models and temperatures.
+Our analysis reveals several notable patterns in the networks generated by LLM agents under different conditions.
+First, models newer than GPT-3.5 exhibit a slightly larger $\hat{\gamma}$ than GPT-3.5.
+This implies that these models display a stronger tendency toward preferential attachment and the formation of hubs.
+Second, as the temperature increases the power-law exponent $\hat{\gamma}$ generally becomes larger.
+This indicates that higher temperatures introduce more variance in node connectivity, leading to degree distributions with heavier tails.
+Third, the environmental context significantly affects the value of $\hat{\gamma}$.
+For example, when the network is framed within a “school” environment, the exponent increases, suggesting a more uniform distribution of connections and fewer highly central nodes.
+Finally, while the prompts we have utilized thus far have provided the model with the complete existing network structure, we also explore an alternative scenario: what happens if agents are supplied solely with the degree of other alternatives, without access to the network’s full structure?
+As detailed in Appendix B, our findings reveal that limiting agents to degree information alone also leads to notable structural differences in the networks that emerge (cf Figure S.9).
+Thus, degree information alone yields more restrictive structures than providing the agents with the full topological information (i.e.
+the neighbors).
+The findings highlight the practical potential of LLMs in modeling complex networks, such as social, economic, or biological systems, by leveraging their ability to simulate preferential attachment and scale-free distributions.
+These models can be used to study real-world phenomena like information diffusion, hub formation, or connectivity patterns under varying conditions.
+Additionally, the sensitivity of network structures to parameters like temperature and context underscores the importance of prompt design in steering outcomes, making LLMs versatile tools for tailored simulations.
 Principle 2: Triadic Closure
-(a) Structure-Based
-(b) Common-Neighbor-Based
+(a) Probability of connecting to top-$k$ for different models, temperatures, and environments
+(b) Marginal transitivity ($D$) and probability of an edge within a community ($\hat{p}$) for different models, temperatures, and environments
+(c) Network instances by GPT-3.5 agents
 Figure 2.
 Results for Principle 2 (triadic closure).
-(a) Figure 2(a) shows the resulting networks created by triadic closure when the intersection of the neighborhoods of the query node and each alternative is provided.
-The node colors correspond to the stochastic block model (SBM) groups to which each node belongs.
-The red edges correspond to the newly created edges between nodes of the red cluster, the blue edges correspond to the newly created edges between nodes of the blue cluster, and the orange edges correspond to edges created between the two clusters.
-The newly formed edges are drawn with a larger line width.
-We observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test).
-(b) Figure 2(b) shows the same networks with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative.
-Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test).
-The error bars correspond to 95% confidence intervals.
+Top: Probability of connecting to top-$k$ nodes (in terms of common neighbors) for varying model (temperature is fixed to 1.0 and environment to baseline), temperature (model fixed to GPT-4 Mini and environment to baseline) and environment (model fixed to GPT-4 Mini and environment temperature to 0.5) for networks generated according to Principle 2 ($n=50$, 10 simulations for each model, environment and temperature).
+Middle: Marginal transitivity ($D$) and probability of an edge within a community ($\hat{p}$) for networks generated according to Principle 2 in different models, temperatures, and environments.
+Bottom: The figure shows the resulting networks created by GPT-4 Mini, according to Principle 2 when the intersection of the neighborhoods of the query node and each alternative is provided.
+The node colors correspond to the groups to which each node belongs.
+The bold edges (red or blue) correspond to the newly created inter-cluster edges, and the orange edges correspond to the new intra-cluster edges.
 The second micro-level principle we examine is triadic closure, which posits that individuals are more likely to form connections with friends of friends, thus creating closed triads in the network.
@@
 This setup mirrors our assumption that nodes within the same cluster are more inclined to connect due to a higher number of shared neighbors.
-In subsequent time steps, we then examine each node $i$, considering the intersection of neighborhoods of $i$’s non-neighbors or the number of mutual friends between $i$ and any non-neighbor $j$ [^5] It’s worth noting that similar outcomes arise when providing neighbors instead of common neighbors.
+In subsequent time steps, we then examine each node $i$, considering the intersection of neighborhoods of $i$’s non-neighbors [^2] Similar outcomes arise when providing neighbors instead of common neighbors..
 We conducted ten simulations with $n=50$ nodes to facilitate clear visualization and ensure statistical significance [^3] Choosing $n=50$ instead of a larger number like $n=200$ aids in visualization and maintains statistical significance..
-To contrast LLM-based network formation, we refer to a random null model, selecting nodes at random for connection.
-For evaluating triadic closure, we utilize two metrics: marginal transitivity and probability of edge formation within the same community.
+On a micro-scale, Figure 2(a) illustrates the probability of connecting to a top-$k$-percentile node as a function of the number of common neighbors.
+The dashed lines represent the results of null models, where connections are chosen randomly; which corresponds to the probability of connecting to a top-$k$ percentile node in terms of the common neighbors being $k/n$.
+Our findings reveal that, across all models, there is a consistently higher probability of forming links with nodes that share more common neighbors.
+Unlike the behavior observed in preferential attachment, temperature does not appear to severely impact this probability.
+This tendency to form links with nodes that have more common neighbors is consistent across various contexts, including school, work, and community environments.
+These results suggest that the triadic closure tendency is a robust phenomenon, persisting across different model families, configurations, and environments.
+Then, for evaluating triadic closure on the network (macroscopic) level, we utilize two metrics: marginal transitivity and probability of edge formation within the same community.
 Marginal transitivity ($D$) represents the change in the ratio of closed triangles to all triads, transitioning from the initial network $G_{T}$ to the SBM-generated $G_{1}$:
@@
 As we investigate under SBM, same community membership indicates more open triads being closed.
-In Figure 2(a), sample networks are displayed, with the upper panel showing networks where the entire structure is provided and the lower panel showing those with only common neighbor numbers provided.
+Marginal transitivity ($D$), presented in Figure 2(b), demonstrates a statistically significant increase across all models, temperatures, and environments, underscoring the robust nature of triadic closure.
+Similarly, the probability of within-community edge formation ($\hat{p}$), shown in Figure 2(b), consistently exceeds $0.5$, reaffirming the tendency of nodes to form new connections within their respective clusters.
+In Figure 2(c), sample networks from GPT-3.5 are displayed, with the upper panel showing networks where the entire structure is provided and the lower panel showing those with only common neighbor numbers provided.
 Nodes are color-coded to indicate their cluster memberships in the SBM, with red and blue edges within clusters and orange edges between clusters.
 Newly formed edges are highlighted with thicker lines.
-Marginal transitivity, presented in the fourth panel of each row, shows a statistically significant increase across all temperatures ($P<0.001$, t-test), and compared to the Random model, the increase is notable ($P<0.001$, t-test).
-Regarding the probability of within-community network formation (last panels in each row), we found that across all temperatures, this probability significantly exceeds that in the random network formation scenario ($P<0.001$, t-test against Random), indicating an LLM tendency towards triadic closure.
-These findings affirm that LLM agents tend to follow triadic closure.
-We investigate the reasons provided by LLMs for their decisions in SI Section 3.777Reasons for decisions based on the number of common neighbors are not provided, as these decisions are solely based on this information.
-The analysis revealed that many decisions were influenced by the number of mutual friends.
+In summary, these findings show that most LLMs exhibit a consistent tendency for triadic closure across various configurations, temperatures, and environments.
+This behavior mirrors human network dynamics, highlighting the models’ ability to simulate realistic social and structural networks and reinforcing their alignment with social principles observed in real-world communities.
 Principle 3: Homophily
+(a) Assortativity and Louvain Modularity with different LLM models and environments
+(b) Network instances generated by GPT-3.5 agents
+Figure 3.
+Results for Principle 3 (Homophily) and Principle 4 (Community structure due to homophily).
+Top: Assortativities and Louvain modularity according to Principle 3 ($n=50$, 5 simulations for each row) in different environments (school, work, community) using different models.
+The statistical significance is $P<0.001$ for all t-tests (comparing with 0).
+Bottom: Network instances and community structure.
+Figure 4. Effect of distractor features (favorite color and lucky number) on homophily.
 Homophily reflects the tendency for nodes with similar characteristics or attributes to form connections and associate with each other.
@@
 The seed network is taken to be the empty graph.
-We run ten simulations for networks with $n=50$ nodes and $\delta=5$. We compare with the Random null model, which chooses $\delta=5$ links at random.
-To measure the presence of homophily, we measure the attribute assortativity coefficient of each of the three features (hobby, color, location).
+We run ten simulations for networks with $n=50$ nodes and $\delta=5$.
+To evaluate homophily, we calculate the attribute assortativity coefficient for each of the features.
 For each property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$ (indexed by $k$ or $l$), its assortativity coefficient $R$ is defined as
@@
 Conversely, a negative assortativity suggests connections primarily occur between dissimilar nodes, indicating heterophily.
-In Figure 3(a), we observe a consistent pattern of positive assortativity across all temperature ranges, which is higher for lower temperatures.
-Moreover, the observed assortativity at all temperatures significantly differs from Random, which has assortativity close to 0 ($P<0.001$, t-test).
-Moreover, we present the probability of each reason being mentioned in SI Section 3.
-We find that shared hobbies appear to be much more important than shared location or favorite color.
-This is consistent with our intuition that, similarly to humans, LLMs care more about hobbies than other superficial attributes when forming social connections.
-In token, we conclude homophily emerges for the networks generated by LLMs.
+From Figure 3(a), we observe that different attributes exhibit varying levels of assortativity.
+First, homophily is present across all LLMs—regardless of the specific model or configuration (e.g., temperature settings), all show positive assortativity for all four attributes.
+This aligns with human societies, where homophily is a primary driver of network formation (McPherson et al., 2001) [^4] As an additional robustness check, we also tested mutual agreement connections.
+In that setting, after a node $j$ is chosen by node $i$, $j$ has to confirm the creation of the link from itself to $i$ ($j\to i$).
+We ran several experiments with different models and temperatures and we found the results not to be affected, namely, the proposed connections were always bilateral..
+Moreover, to test the effect of the features on homophily as indicated by the assortativity coefficient for each attribute, we introduce a distractor feature, which corresponds to a lucky number that is randomly chosen between 0 and 9.
+We repeat the simulations for all models and measure the effect of each feature on Figure 4.
+We show that lucky numbers consistently show lower assortativity coefficients, indicating they are less considered when forming homophilous connections.
+This is consistent with our prior expectation that humans typically do not prioritize shared lucky numbers when establishing relationships.
+Surprisingly, even though the lucky number does not seem to impact homophily much, the favorite color exhibits a similar level of homophily as hobbies.
+One might expect that hobbies, being substantive interests, would have a stronger influence on social connections than favorite colors, which are more arbitrary preferences.
+However, this finding aligns with the social identity theory and the minimal group paradigm (Tajfel et al., 1971).
+According to this paradigm, even minimal and arbitrary group distinctions – such as a preference for certain colors – can lead to in-group favoritism and influence social connections.
+This suggests that LLM agents, akin to humans, may form connections based on even trivial shared attributes, reflecting inherent tendencies toward group formation based on minimal commonalities.
+All in all, LLMs can capture and reproduce subtle human social behaviors, not just linguistic patterns.
+This underscores their potential as powerful tools for social simulation.
+However, these findings may also raise important considerations regarding bias, fairness, and the ethical design of AI systems (cf.
+Discussion Section).
 Macro-Level Principles
-(a) Assortativity Coefficients
-(b) Network Instances
-(c) Communities Statistics
-Figure 3.
-Results for Principle 3 (homophily).
-(a) Figure 3(a) shows the assortativity coefficients for each network attribute created based on Principle 3.
-Each node can form at most $\delta=5$ links with the other nodes based on their attributes.
-We observe that for all temperatures, we have positive assortativity ($P<0.001$, t-test with 0), which indicates homophily.
-The assortativities are significantly different for all temperatures from random or heterophilous ($P<0.001$, t-test comparing with Random).
-(b) Networks created based on Principle 3.
-Each node can form at most $\delta=5$ links with the other nodes based on their attributes.
-The edge widths correspond to the number of common attributes shared between the nodes.
-The node colors correspond to the communities found by running the Louvain algorithm.
-(c) Louvain modularity, number of communities, and community sizes of networks created based on Principle 3.
-Lower temperatures yield higher modularity ($P<0.001$, t-test) and fewer and larger communities ($P<0.001$, t-test).
-For smaller temperatures, the communities are more “well-separated”.
-The error bars correspond to 95% confidence intervals.
 Principle 4: Community Structure
@@
 Identifying community structures is crucial for understanding the overall dynamics of a network, as it reveals patterns of relationships and interactions that might not be apparent at the global level.
+[^5] As an example, we present only the results from GPT-3.5 for Principle 4 (Community Structure) and Principle 5 (Small-World).
 Both triadic closure and homophily contribute to the formation of community structures.
@@
 Here, $x_{i}^{(k)}$ and $x_{j}^{(k)}$ correspond to the $k$-th features of $x_{i}$ and $x_{j}$, respectively.
-In Figure 3(b), various colors represent the communities identified by the Louvain algorithm at different temperatures.
+In Figure 3(b), various colors represent the communities identified by the Louvain algorithm at different temperatures for GPT-3.5.
 Notably, communities appear more distinct at lower temperatures, likely due to reduced randomness in decision-making at these temperatures.
-Additionally, Figure 3(c) presents the distribution of Louvain modularity values across simulations, indicating consistent community structure with positive modularity at all temperatures, confirmed by a t-test against a modularity of $Q=0$ for a random graph ($P<0.001$).
-Moreover, lower temperatures lead to fewer, larger communities as per the Louvain algorithm’s findings ($P<0.001$, t-test).
+Figure 3(a) presents the distribution of Louvain modularity values across simulations accross different LLM models and different environments, indicating consistent community structure with positive modularity at all temperatures, confirmed by a t-test against a modularity of $Q=0$ for a random graph ($P<0.001$).
 Our results demonstrate that community structures manifest in networks generated by LLMs, driven by both triadic closure and homophily.
@@
 (f) $\beta=0.75$
+(g) Regression plot for different models and environments for $\beta=0.25$ and $k=5$.
 Figure 6.
 Fitted results for Principle 5 (small world).
-Top: Regression plots relating average shortest path length ($L$) and average clustering coefficient ($C$) with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$.
-Value $a$ in legends represents the slope of the regression lines.
-Bottom: Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with the estimated rewiring probability $\hat{\beta}$.
+Top (a-c): Regression plots relating average shortest path length ($L$) and average clustering coefficient ($C$) with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$ for GPT-3.5.
+The value $a$ in legends represents the effect size (slope of the regression lines).
+Middle (d-f): Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with the estimated rewiring probability $\hat{\beta}$ for GPT-3.5 agents.
 We report the $P$-values of the t-test comparing the average shortest path length of the LLM-generated networks and the average shortest path length of the Watts-Strogatz graphs with rewiring probability $\hat{\beta}$.
-*: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.
+Bottom (g): Regression plot for the relation $L\sim\log(n)$ for different LLM models and environments (school, work, community) for $\beta=0.25$ and $k=5$.
+The legend shows the effect size ($a$) and the $P$-value.
+(*: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.)
 The small-world phenomenon is characterized by networks where nodes are interconnected in tight clusters, yet the average distance between any two nodes remains relatively short, typically scaling logarithmically with the network size (Watts and Strogatz, 1998; Kleinberg, 2000).
@@
 For this analysis, we considered values of $\beta$ set at 0.25, 0.5, and 0.75, with a fixed $k=5$ to serve as a consistent parameter.
-Visual representations and average clustering coefficients are presented in Figures 4(a), 4(b) and 4(c).
+Visual representations and average clustering coefficients are presented in Figures 5(a), 5(b) and 5(c).
 However, when directly compared with the Watts-Strogatz model, the networks generated by the LLMs do not precisely replicate the characteristics of Watts-Strogatz networks for the corresponding the rewiring probabilities ($\beta$).
-As illustrated in right panels of Figures 4(a), 4(b) and 4(c), we find weak evidence that the LLM-generated networks share the same average shortest path length as the Watts-Strogatz model for the rewiring probabilities ($\beta$) of 0.25, 0.5, and 0.75, with $P<0.1$ from a t-test comparing the average shortest path lengths.
+As illustrated in right panels of Figures 5(a), 5(b) and 5(c), we find weak evidence that the LLM-generated networks share the same average shortest path length as the Watts-Strogatz model for the rewiring probabilities ($\beta$) of 0.25, 0.5, and 0.75, with $P<0.1$ from a t-test comparing the average shortest path lengths.
 Additionally, LLM-generated networks exhibit a larger average clustering coefficient than those of the Watts-Strogatz model for the same rewiring probabilities $\beta$, also with a significance level of $P<0.1$ in the t-test comparisons.
@@
 [^8] We conducted a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks.
-In Figures 5(d), 5(e) and 5(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$ and each temperature.
+In Figures 6(d), 6(e) and 6(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$ and each temperature.
 Here $P$-values result from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$.
 These results show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings.
+Finally, as Figure 6(g) shows, the relation $L\sim\log(n)$ holds for different LLM models and environments.
 In conclusion, our analysis demonstrates that LLM-generated networks exhibit key small-world properties, with logarithmic scaling of average shortest path lengths and inverse logarithmic scaling of average clustering coefficients.
 While these networks do not perfectly align with the Watts-Strogatz model, they exhibit similar structural characteristics.
-Decisions on Real-World Networks
-To investigate the behavior of LLMs in real-world network formation contexts, we employ the Facebook100 dataset (Traud et al., 2012) and a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972).
-This dataset comprises “friendship” networks from one hundred American colleges and universities, captured at a specific moment in time from Facebook’s online social network.
-We model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives (see Methods and Materials).
-We use datasets that refer to three US colleges – Caltech36, Swarthmore42, and UChicago30.
-They represent different sizes of networks: $n=769$, $n=1,659$, and $n=6,591$ nodes, respectively.
-For the Caltech36 ($n=769$) and Swarthmore42 ($n=1,659$) datasets, we consider all of the network nodes to be nodes that have to make a choice.
-For the UChicago30 dataset, we consider a randomly sampled subset of 2,000 nodes because of the limited context window of GPT models.
-We set the size of the alternative sets $A_{t}$ to 20, and thus, the accuracy of random guessing is 5%.
-Figure 6.
-Regression results for real-world network experiments.
-Multinomial logit coefficients for three networks from the Facebook100 dataset.
-The standard errors of the estimates are shown with error bars.
-The full regression tables can be found in SI Section 1.
-All coefficients are statistically significant, where the null hypothesis corresponds to the respective parameter being equal to 0.
+Decisions on Real-World Networks with Heterogeneous Agents
+We investigate the behavior of LLMs in real-world network formation contexts with four datasets in two differing real-world domains.
+Despite the significant advancements in social network analysis over recent years, the availability of fully complete and comprehensive network datasets remains exceptionally rare (Yuan et al., 2018).
+We employ three datasets from the Facebook100 collection (Traud et al., 2012) and the telecommunication (Andorra) and the employment (MobileD) datasets from (Yuan et al., 2018).
+The Facebook100 data correspond to “friendship” networks from one hundred American colleges and universities, captured at a specific moment from Facebook’s online social network.
+The Andorra dataset contains nationwide call records in Andorra from July 2015 to June 2016, where calls correspond to mutual calls between Andorran residents containing information about the caller’s and the callee’s location, phone type, and usage.
+Finally, the MobileD dataset corresponds to a company network where relations correspond to call or text communication, and each employee is either a manager or a subordinate.
+For all network datasets, the agents have heterogeneous profiles (i.e., profiles with different features) whose statistics (degree distribution, clustering coefficient distribution, assortativity) we report in Figure 7.
+To infer the models’ tendencies, we employ a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972).
+Specifically, we model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives (see Methods and Materials).
+Number of Samples and Alternatives
+For the three datasets from Facebook100 are Caltech36 ($n=769$) Swarthmore42 ($n=1,659$), and UChicago30 ($n=6,591$), we set the number of alternatives to be $A=15$ and randomly sampled from the existing network.
+For the UChicago30 dataset, we consider a randomly sampled subset of $N=2,000$ nodes because of the limited context window of the LLM models.
+For Andorra ($n=32,812$) and MobileD ($n=1,982$), we set the number of alternatives to $A=5$ and consider a randomly sampled subset of $N=1,000$ nodes.
+Figure 7. Distributions of real-world datasets analyzed in our study, including degree, clustering coefficients, and the assortativities of the attributes included in the datasets.
+Model Preferential Attachment ($\hat{\theta}_{\mathrm{PA}}$) Homophily ($\hat{\theta}_{\mathrm{H}}$) Triadic Closure ($\hat{\theta}_{\mathrm{TC}}$) Log Likelihood AIC
+Caltech36 ($n=769$ nodes, $m=33,312$ edges, $N=769$ samples, $A=15$ alternatives each)
+GPT-3.5 0.20*** (0.002) 0.65*** (0.005) -0.06 (0.006) -2,088.21 4,184.41
+GPT-4o Mini 0.34*** (0.006) 2.13*** (0.03) 0.44*** (0.02) -1,201.27 2,410.55
+GPT-4 (gpt-4-1106-preview) 0.41*** (0.01) 1.95*** (0.02) 0.59*** (0.01) -1,377.47 2,762.94
+Claude 3.5 Sonnet 0.46*** (0.005) 0.55*** (0.01) 0.55*** (0.007) -1,748.19 3,504.38
+Llama 3 70b Instruct 0.28*** (0.006) 2.43*** (0.02) 0.84*** (0.01) -809.57 1,627.15
+Swarthmore42 ($n=1,659$ nodes, $m=122,100$ edges, $N=1,659$ samples, $A=15$ alternatives each)
+GPT-3.5 0.19*** (0.008) 0.47*** (0.01) 0.00 (0.009) -4,484.45 8,976.90
+GPT-4o Mini 0.27*** (0.21) 2.22*** (0.78) 0.57*** (0.43) -1,899.09 3,806.19
+GPT-4 (gpt-4-1106-preview) 0.18*** (0.003) 1.62*** (0.006) 0.65*** (0.002) -2,838.33 5,684.66
+Claude 3.5 Sonnet 0.36*** (0.002) 0.75*** (0.006) 0.55*** (0.004) -3,563.02 7,134.03
+Llama 3 70b Instruct 0.39*** (0.003) 2.31*** (0.005) 0.62*** (0.004) -1,820.26 3,648.52
+UChicago30 ($n=6,951$ nodes, $m=416,206$ edges, $N=2,000$ samples, $A=15$ alternatives each)
+GPT-3.5 0.22*** (0.001) 0.48*** (0.004) -0.02 (0.0005) -8,157.38 16,322.77
+GPT-4o Mini 0.27*** (0.005) 2.22*** (0.019) 0.57*** (0.011) -1,899.09 3,806.19
+GPT-4 (gpt-4-1106-preview) 0.23*** (0.001) 2.00*** (0.005) 0.41*** (0.002) -3,444.33 6,896.67
+Claude 3.5 Sonnet 0.43*** (0.003) 0.78*** (0.005) 0.39*** (0.002) -6,604.77 13,217.54
+Llama 3 70b Instruct 0.43*** (0.007) 2.57*** (0.014) 0.32*** (0.005) -3,689.00 7,386.00
+Andorra ($n=32,812$ nodes, $m=513,931$ edges, $N=1,000$ samples, $A=5$ alternatives each)
+GPT-3.5 0.53*** (0.001) 0.21* (0.01) -0.24*** (0.002) -1,712.91 3,433.83
+GPT-4o Mini 0.54*** (0.004) 3.47*** (0.06) -0.09* (0.01) -1,002.11 2,012.22
+GPT-4 (gpt-4-1106-preview) 0.21 (0.20) 3.45*** (1.16) -0.22 (0.25) -1271.20 2622.30
+Claude 3.5 Sonnet 0.54*** (0.003) 1.94*** (0.009) -0.15*** (0.003) -1,541.77 3,091.55
+Llama 3 70b Instruct 0.38*** (0.003) 3.92*** (0.02) -0.04 (0.01) -985.95 1,979.91
+MobileD ($n=1,982$ nodes, $m=25,470$ edges, $N=1,000$ samples, $A=5$ alternatives each)
+GPT-3.5 1.06*** (0.003) -0.94*** (0.009) -0.02 (0.001) -1,663.42 3,334.84
+GPT-4o Mini 1.38*** (0.02) -0.85*** (0.02) 0.87*** (0.01) -880.39 1,768.78
+GPT-4 (gpt-4-1106-preview) 0.42*** (0.01) -1.84*** (0.009) 0.94*** (0.003) -1,321.01 2,650.02
+Claude 3.5 Sonnet 0.71*** (0.009) -2.44*** (0.02) 1.13*** (0.005) -1,197.92 2,403.83
+Llama 3 70b Instruct 1.04*** (0.005) -0.36** (0.01) 0.71*** (0.002) -1,269.42 2,546.83
+Note: *: $P<0.05$, **: $P<0.01$, ***: $P<0.001$
+Table 1.
+Effect sizes for real-world networks from Facebook100 (Traud et al., 2012) and Andorra dataset (Yuan et al., 2018) for several LLMs for temperature set to 0.5.
+An earlier version of this paper included experiments with gpt-4-1106-preview beyond gpt-4o-mini.
+We have also added these experiments for completeness.
+Regression Coefficients
 We regress network formation decisions on standardized scores reflecting the three micro-level principles.
-We present the regression results in Figure 6.111111More detailed results can be found in Table S1 in SI Section 1.
-Our analysis reveals that the degree, the number of common neighbors, and the number of shared interests are all statistically significant factors influencing the choices of LLMs ($P<0.001$) across all datasets.
-Furthermore, we find that LLMs place the greatest emphasis on the number of common attributes ($\hat{\theta}_{\mathrm{H}}$), followed by the number of common neighbors ($\hat{\theta}_{\mathrm{TC}}$), and then the degree ($\hat{\theta}_{\mathrm{PA}}$).
-These results suggest that LLMs exhibit a strong preference for homophily, followed by a tendency towards triadic closure, and finally, preferential attachment.
-Specifically, the relationship $\hat{\theta}_{\mathrm{H}}>\hat{\theta}_{\mathrm{TC}}>\hat{\theta}_{\mathrm{PA}}$ holds across all examined datasets and temperature settings.
-Further insights are provided in SI Section 4, where we provide the reasoning analysis.
-The analysis points to a marked emphasis on homophily, indicating that nodes with similar characteristics—such as student/faculty status, high school, and graduation year—are more likely to connect.
-In contrast, triadic closure and preferential attachment have a relatively lesser impact.
-Our study also includes a comparison of LLMs’ predictive capabilities against random guessing in a link prediction framework.
-We assess the percentage increase in accuracy, defining accuracy as the probability of selecting an existing neighbor from a set of alternatives, as opposed to a random selection from 20 possible options.
-As detailed in SI Section 1, LLMs significantly outperform random guessing in accuracy.
-Furthermore, we investigate whether LLMs reinforce community structure post-network formation.
-We run the Louvain algorithm 10 times for both the initial graph $G_{1}$ and the final graph $G_{T}$ and perform a t-test between the resulting modularity scores (see SI Section 1).
-We observe that the modularity increases in all cases and is statistically significant.
-This suggests that LLMs make choices that naturally yield a reinforcement of the community structure.
-Additionally, to explore the “small-world” phenomenon, we measure the percent change in average shortest path length between the pre- and post-decision networks (detailed in SI Section 1).
-Our findings reveal a minor decrease in average path length in most cases, suggesting a slight strengthening of the small-world phenomenon.
+We present the regression results in Table 1.999More detailed results can be found in Table S.2 in Appendix A.
+First, we observe a dominant effect of homophily across all datasets and models.
+The coefficients for homophily ($\hat{\theta}_{\mathrm{H}}$) are consistently the largest and highly significant ($P<0.05$) in almost all cases.
+For instance, in the Caltech36 dataset, the homophily coefficients for GPT-3.5, GPT-4, and Llama 3 70b Instruct are 0.65, 1.95, and 2.43, respectively ($P<0.001$).
+The emphasis on homophily suggests that LLMs, much like humans, prioritize forming connections based on shared characteristics.
+Second, we find that preferential attachment plays a secondary role in the network formation decisions of LLMs.
+While the coefficients for preferential attachment ($\hat{\theta}_{\mathrm{PA}}$) are generally positive and statistically significant across most models and datasets, they are notably smaller than those for homophily.
+For example, in the Swarthmore42 dataset, GPT-3.5 and Llama 3 70b Instruct have preferential attachment coefficients of 0.19 and 0.39, respectively ($P<0.001$).
+This suggests that while LLMs do consider the degree of potential connection nodes—favoring connections to well-connected nodes—the influence of this factor is less evident compared to homophily.
+Finally, the influence of triadic closure appears to vary across different datasets and models.
+In most cases, the coefficients for triadic closure ($\hat{\theta}_{\mathrm{TC}}$) are positive and significant, indicating that LLMs consider the number of mutual connections when forming new links.
+However, in some instances, such as with GPT-3.5 on the Andorra dataset, the triadic closure coefficient is negative ($-0.24$) and significant, suggesting a structure-dependent role of this principle as shown by the low clustering coefficient of the network, which is dominated by preferential attachment, as also shown in Figure 7.
+This variability implies that while triadic closure is a factor in LLMs’ decision-making, its impact may be influenced by the specific characteristics of the dataset or the model used.
 In summary, our analysis demonstrates that while homophily, triadic closure, and preferential attachment are integral to the network formation behaviors of LLMs, homophily is the dominant factor.
-Furthermore, both community structure and the small-world phenomenon are reinforced after network formation decisions of LLMs.
 Discussion
-Our study conducted a comprehensive evaluation of LLMs’ network formation preferences, encompassing both micro-level network principles such as preferential attachment, triadic closure, and homophily and macro-level network principles like community structure and the small-world phenomenon.
-Our findings indicate that networks generated by multiple LLMs do exhibit these properties, particularly when primed with network statistics such as the number of mutual friends or the degree of alternatives.
-Furthermore, through discrete choice modeling, we explored the emergence of these properties in real-world networks.
-Our real-world network results reveal that the agents’ selections are predominantly driven by homophily, followed by triadic closure, and then preferential attachment.
-Additionally, we observed a strengthening of community structure in real-world network experiments.
-Our study enhances our understanding of the behavior of multi-LLMs.
-For example, our findings reveal different strengths of network formation properties, suggesting that when LLMs are employed to coordinate social networks in social or work environments, they may exhibit behaviors similar to humans.
-This insight contributes to developing socially aware LLMs that align with human preferences, with potential applications ranging from chatbots and virtual assistants to collaborative systems and social media platforms.
-Furthermore, our study may pave the way for innovative applications of LLMs in network science and computational social science research.
-Given that LLMs adhere to fundamental network principles and exhibit behaviors akin to human network formation, they hold significant potential for use in simulations of social systems and agent-based modeling.
-As LLM agents gain increasing ability to exhibit more complex and nuanced behaviors, such as negotiation, collaboration, and competition, this presents opportunities to replicate networked environments like marketplaces or organizational structures.
-In these settings, LLMs could effectively model decision-making processes and interaction dynamics.
-Looking ahead, we propose three promising directions for future research.
-First, we propose investigating LLM behavior in more complex interactions, such as simulated dialogs or specific environments like companies, schools, or parties.
-By examining LLMs in these contexts, we can better understand their network formation preferences and how they adapt to different social dynamics.
-This research could shed light on the nuances of LLM behavior in realistic settings and inform their integration into diverse social environments.
-Second, we suggest exploring the application of our findings in specific settings, particularly in organizational contexts.
-For example, considering the potential use of LLMs to assist HR professionals, our study’s insights could guide the alignment of LLMs with organizational goals.
-By leveraging network information, LLMs could help HR departments identify their firms’ best fit or talent, optimizing recruitment and talent management processes.
-Finally, we propose using our methods to create realistic synthetic networks.
-These synthetic datasets can serve as benchmarks for evaluating graph learning methods, thus addressing the scarcity of existing graph benchmarks.
-By adjusting parameters like temperature, we can generate diverse networks to test graph neural network performance under various conditions.
-In addition, our approach can generate artificial data resembling real-world data while adhering to privacy regulations.
-This has the potential to expedite research in graph learning and trustworthy machine learning by providing realistic network datasets generated by LLMs.
-1. Methods
+Summary of Findings and Broader Impact
+In this study, we conducted a comprehensive evaluation of LLMs’ network formation preferences, examining both micro-level network principles—such as preferential attachment, triadic closure, and homophily—and macro-level network properties like community structure and the small-world phenomenon.
+Our findings indicate that networks generated by multiple LLMs exhibit these properties, particularly when the models are primed with network statistics like the number of mutual friends or the degrees of potential connections.
+Furthermore, using discrete choice modeling, we explored the emergence of these properties in simulations based on real-world networks.
+Our results reveal that the LLM agents’ selections are predominantly driven by homophily, followed by triadic closure and preferential attachment.
+On the one hand, our study enhances our understanding of how multiple LLMs behave in networked settings.
+Specifically, our findings reveal varying strengths in network formation properties among LLMs, suggesting that when these models are employed to coordinate social networks in social or work environments, they may exhibit human-like behaviors.
+This has important implications for applications like agent-based modeling, where realistic simulation of human behavior is crucial.
+Traditionally, agent-based models rely on simplified rules or heuristics to represent individual behaviors, which may not capture the complexity of human decision-making.
+By incorporating LLMs as agents, we can simulate more nuanced and context-aware interactions that closely resemble human social behavior, without the need to rigidly specify decision rules or heuristics.
+Limitations and Boundary Conditions
+On the other hand, our results suggest that we should exercise caution when leveraging LLMs in networking scenarios.
+The model family, configuration, and prompts can subtly affect the models’ behavior, resulting in qualitatively similar but quantitatively different outcomes.
+For example, newer models such as GPT-4 and Claude 3.5 exhibit stronger biases compared to prior models such as GPT-3.5.
+For instance, in the preferential attachment principle, newer models such as GPT-4 and Claude 3.5 have stronger biases – i.e., connect to highest-degree-nodes yielding star-like networks – compared to GPT-3.5 and LLama 3, which had weaker biases – i.e., connect to high-degree-nodes.
+Similar results can be found in homophily, as larger biases towards homophily, and triadic closure.
+Thus, even though LLMs exhibit these principles, we should be cautious of such biases when designing simulations.
+Similarly, in our experiments with real-world data, we find an interesting phase transition from homophily to heterophily: In the Facebook100 data, the LLMs generally exhibited positive biases toward homophily ($\hat{\theta}_{H}>0$).
+However, in employment networks, agents were either managers or subordinates, and we discovered that LLM agents were heterophilous in such a case ($\hat{\theta}_{H}<0$), which aligns with career advancement dynamics (i.e.
+employees want to form links with managers because of better career prospects).
+The above underscores the need for researchers to provide oversight and ensure that LLM behaviors align with human expectations when employing them in scientific research methods, such as agent-based modeling and even prototypical human subject research with LLMs.
+Thus, although we find that LLMs resemble human network formation behaviors, we should consider whether these models should exhibit such behaviors when serving as assistants to humans in work and social lives.
+Biases like homophily, triadic closure, or preferential attachment may lead to network structures that overemphasize certain individuals or fragment information flow.
+Humans form networks with communities and central nodes partly due to limited social capacities; however, LLMs do not share these limitations.
+Therefore, when used as social assistants, LLMs may not necessarily need to mirror human networking behaviors and could be personalized to promote more equitable and efficient information dissemination.
+Our study highlights the need for more deliberate efforts to align LLM behavior in this domain.
+Future Research Directions
+Looking ahead, we identify three promising directions for future research: First, investigating LLM behavior in more complex interactions, such as simulated dialogues, could provide deeper insights.
+By examining the specific dialogues of LLMs during interactions, we can better understand their network formation preferences and how they adapt to different social dynamics.
+Second, we could explore how LLMs can be integrated into real-world settings, such as social media platforms.
+LLM-assisted bots might be employed to facilitate interactions, break echo chambers, and moderate democratic discussions (see, for example, (Papachristou et al., 2023)).
+Finally, we can use our methods to create realistic synthetic networks.
+These synthetic datasets can serve as benchmarks for evaluating graph learning methods, thereby addressing the scarcity of existing graph benchmarks (cf.
+(Palowitch et al., 2022)).
+By adjusting parameters like temperature and environmental settings or selecting different models, we can generate diverse networks to test graph neural network performance under various conditions.
+Importantly, our approach allows for the generation of artificial data that resembles real-world data while adhering to privacy regulations.
 Methods and Materials
+# Task You are located in a school.
+Your task is to select a set of people to be friends with.
+# Profile Your profile is given below after chevrons: <PROFILE> { "name" : "Person 0", "favorite subject" : "Chemistry", "neighbors" : ["Person 3", "Person 432", "Person 4", "Person 3", "Person 32"] } </PROFILE> # Candidate Profiles The candidate profiles to be friends with are given below after chevrons: <PROFILES> [ { "name" : "Person 1", "favorite subject" : "Mathematics", "neighbors" : ["Person 3", "Person 4", "Person 23", "Person 65"] }, { "name" : "Person 33", "favorite subject" : "History", "neighbors" : ["Person 342", "Person 2", "Person 12"] }, ...
+] </PROFILES> # Output The output should be given a list of JSON objects with the following structure [ {{ "name" : name of the person you selected, "reason" : reason for selecting the person }}, ...
+] # Notes - The output must be a list of JSON objects ranked in the order of preference.
+- You can make at most 1 selection.
+ALGORITHM 1 Example prompt regarding social network data.
 # Task Your task is to select a set of people to be friends with.
@@
 - You can make at most $\delta$ selections.
-ALGORITHM 1 Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$.
+ALGORITHM 2 General Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$.
 Experimental Procedure
@@
 Subsequently, we utilized real-world networks to determine the factors that are most heavily weighted by LLMs.
+Network Formation Process
 Our experiments span a time series of $T$ steps, with a sequence of network structures denoted as $G_{1},G_{2},\dots,G_{T}$ with vertex sets $V_{1},\dots,V_{T}$.
@@
 We employ multiple temperatures to account for the variability in response generation by LLM systems, which is also observed in classical statistical models of network formation (Jackson, 2008).
-Our study conducts experiments using three temperatures: 0.5, 1.0, and 1.5.
+Our study conducts experiments using three temperatures for all models except Claude 3.5: 0.5, 1.0, and 1.5.
+For Claude 3.5 the temperature range is between 0 and 1, and we run experiments with two temperatures: 0.5, and 1.0.
 Moreover, the model is tasked with outputting a JSON object indicating the node chosen for link formation and the rationale behind the choice.
 This approach is adopted because LLMs have demonstrated proficiency in processing code-like structures, such as HTML and JSON.
-Summarizing the Reasoning of the LLM Agents
-In addition to analyzing the innate preferences of LLM agents, we employ an interview-like approach to probe the rationale behind their decisions.
-Following the LLMs’ decisions on network formation, we inquire about their reasoning for making those choices.
-We then aggregate all the provided rationales using GPT models (GPT-3.5 for the synthetic networks and GPT-4 for the real-world networks) by asking it to identify the common themes, resulting in the creation of multiple categories.
-To determine the category to which each LLM’s reasoning belongs, we subsequently ask the GPT models to classify each rationale into the appropriate category.
+Feature Representations for Prompts
+Below, we give examples of the features used in the prompt presented in Algorithm 2. The features are formatted as a list of JSON objects which are provided to the prompt.
+Principle 1: Preferential Attachment. We have the following features:
+"name" : 0,
+"neighbors" : [5, 7, 1, 6]
+Principle 2: Triadic Closure. We have the following features:
+"name" : 0,
+"common_neighbors" : [5, 7, 1, 6]
+Principle 3: Homophily. We have the following features:
+"name" : 0,
+"favorite_color" : "red",
+"hobby" : "hiking",
+"location" : "Boston"
+Principle 5: Small-World. We have the following features:
+"name" : 0,
+"neighbors" : [5, 7, 1, 6]
+Real-World Data. We have the following features:
+"name" : 0,
+"status" : "student",
+"major" : 10,
+"second major" : 93,
+"accommodation" : "house",
+"high_school" : 5,
+"graduation_year" : 2008
+We note that the initial Facebook100 dataset included gender information as a feature.
+We chose not to include gender as one of the features as it has been shown that language models exhibit gender bias (Vig et al., 2020; Kotek et al., 2023; Bordia and Bowman, 2019).
+An example of the prompt using real-world social network data is given at Algorithm 1.
+Robustness Checks
+We tried the following LLM models:
+  • GPT-3.5 (gpt-3.5-turbo)
+  • GPT-4o Mini (gpt-4o-mini)
+  • Llama 3 (llama-3-70b-instruct)
+  • Claude 3.5 Sonnet (claude-3-5-sonnet-20240620).
+For each of the models except Claude 3.5 we used three temperatures: 0.5, 1.0, 1.5.
+For the Claude 3.5 model we used temperatures 0.5 and 1.0 (since the model does not allow temperatures above 1.0).
+Finally, we experimented with different environmental prompts (e.g., friendship, collaboration, community) to test prompt sensitivity.
 Details for Small-World Experiments
-The algorithm of generating LLM networks is described as follows:
+The algorithm for the altered Watts-Strogatz model is described as follows:
   1. (1)
@@
 The likelihood function is convex, and we optimize it with the L-BFGS-B method (Liu and Nocedal, 1989).
-The standard errors of the coefficients are approximated as $\sqrt{-H^{-1}}$ where $H$ is the Hessian matrix of the log-likelihood at $(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})$ (cf.
+The standard errors of the coefficients are approximated as $\sqrt{-H^{-1}/N}$ where $H$ is the Hessian matrix of the log-likelihood at $(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})$ and $N$ is the number of data points (cf.
 (Overgoor et al., 2019; Train, 2009)).
 Data and Code Availability
-Data and code are available at GitHub: https://github.com/papachristoumarios/llm-network-formation.
+Data and code are available on GitHub at the following link:
+The real-world social network data have been taken from the sources of (Traud et al., 2012) and (Yuan et al., 2018).
 Acknowledgements.
@@
 is supported by a scholarship from the Onassis Foundation (Scholarship ID: F ZT 056-1/2023-2024).
-Appendix A Regression Table
-In Table S.1, we report the regression coefficient for the regression in the real-world network data.
+Appendix A Full Regression Table for GPT-4 (gpt-4-1106-preview) and the Facebook100 Data
+In Table S.2, we report the regression coefficient for the regression in the real-world network data for all temperatures and GPT-4 (gpt-4-1106-preview).
 The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 1.C (also shown in Figure 5) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0).
@@
 Table S.2.
-Multinomial logit coefficients for three networks from the Facebook100 dataset.
+Multinomial logit coefficients for three networks from the Facebook100 dataset and GPT-4 (gpt-4-1106-preview).
 The standard errors of the estimates are shown in parentheses.
@@
 We also report the modularity change (t-statistic) of the whole graph inside brackets.
-Appendix B Analytical Regression Tables
+A.1. Analytical Regression Tables for GPT-4 (gpt-4-1106-preview)
 We present the regression tables for all the combinations of coefficients for each of the three real-world network datasets.
@@
 Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
-Table S.2. Multinomial logit coefficients for Caltech36. The standard errors of the estimates are shown in parentheses.
+Table S.3. Multinomial logit coefficients for Caltech36 and GPT-4 (gpt-4-1106-preview). The standard errors of the estimates are shown in parentheses.
 Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ Log Likelihood AIC
@@
 Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
-Table S.3. Multinomial logit coefficients for Swarthmore42. The standard errors of the estimates are shown in parentheses.
+Table S.4. Multinomial logit coefficients for Swarthmore42 and GPT-4 (gpt-4-1106-preview). The standard errors of the estimates are shown in parentheses.
 Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ Log Likelihood AIC
@@
 Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
-Table S.4. Multinomial logit coefficients for UChicago30. The standard errors of the estimates are shown in parentheses.
-Appendix C LLM Reasoning Summarization
-In addition to studying LLM agents’ innate preferences, we use an interview-style approach to understand why they make certain decisions.
-After LLMs form networks, we ask them about their reasoning.
-We then use GPT models (GPT-3.5 for Section 1.A.
-and GPT-4 for Section 1.C) to group their explanations into common themes, creating various categories.
-To classify each LLM’s reasoning, we have the GPT models assign them to the relevant category.
-C.1. Micro-level Principles
-In Figure S.7, we show the reasoning of the LLM agents from the networks of Section 1.A.
-We observe that for Figure 7(a), the well-connectedness reason becomes less important as the temperature increases while having more mutual friends becomes more important as the temperature increases.
-This means that the agents emphasize triadic closure in higher temperatures and decrease emphasis on preferential attachment.
-For Figure 7(b), we observe an opposite pattern in the LLM reasoning.
-A possible explanation for this asymmetry is the fact that different types of features are provided, i.e., in the case of Figure 7(a) the neighbors of each node are provided, whereas in the case of Figure 7(b) the common neighbors with the query node $i_{t}$ are provided.
-Finally, Figure 7(c) contains the reasons for Figure 3 based on having a common hobby, a common color, or a common location.
-We observe that the LLM agents make choices primarily based on sharing the same hobby and then based on a common location or a common color (on an almost equal basis).
-(a) Reasons for Figure 1(a)
-(b) Reasons for Figure 2(a)
-(c) Reasons for Figure 3(b)
+Table S.5. Multinomial logit coefficients for UChicago30 and GPT-4 (gpt-4-1106-preview). The standard errors of the estimates are shown in parentheses.
+Appendix B Network Evolution and Omitted Simulations
+Here we depict the evolution of the networks generated by the LLM agents, as well as omitted simulations.
+B.1. Principle 1: Preferential Attachment
+Network Evolution
+We plot the evolution of the LLM-based preferential attachment networks at three timesteps, together with the degree distribution alongside the degree distribution of a BA graph with the same number of nodes.
+We observe that for the temperature being 0.5 we have a core-periphery-like formation which diverges from the BA model, whereas for the temperature being 1.5 the network has the same degree distribution as the BA model.
+Figure S.8. Dynamic evolution of networks created based on Principle 1.
+Simulations with Degree Information
+In Figure S.9 we provide the results with degree-information only.
+We observe that the agents form connections around high-degree nodes only (see Figure S.9).
+The same result (star-like networks) holds for the other LLM models and temperatures.
 Figure S.9.
-Distribution of reasoning across categories for the networks of Figures 1(a), 2(a), and 3.
-The classification uses GPT-3.5 on 20 randomly sampled reasons repeated over five times (micro-averaged frequencies).
-The error bars correspond to 95% confidence intervals.
-C.2. Real-World Data
-We also present the LLM reasoning for real-world networks.
-The reasons have been set so that there is a one-to-one correspondence with the regression coefficients.
-Specifically, the Mutual Friends reason corresponds to $\hat{\theta}_{\mathrm{TC}}$, the Number of friends reason to $\hat{\theta}_{\mathrm{PA}}$, and the Similar attributes reason corresponds to $\hat{\theta}_{\mathrm{H}}$ respectively.
-Consistently with the results reported in Table S.1, the agents’ reasoning is more geared towards homophily (same attributes), then triadic closure (mutual friends), and finally preferential attachment (number of friends).
-These results are consistent with the regression.
-Figure S.8.
-Distribution of reasoning across categories for the real-world networks.
-The classification is done using GPT-4 on a set of 20 randomly sampled reasons repeated over 5 times.
-In agreement with Table S.1, the multi-LLMs put more weight on homophily, then on triadic closure, and then on preferential attachment.
-The error bars correspond to 95% confidence intervals.
-Appendix D Network Evolution
-Here we depict the evolution of the networks generated by the LLM agents.
-D.1. Principle 1: Preferential Attachment
-We plot the evolution of the LLM-based preferential attachment networks at three timesteps, together with the degree distribution alongside the degree distribution of a BA graph with the same number of nodes.
-We observe that for the temperature being 0.5 we have a core-periphery like formation which diverges from the BA model, whereas for the temperature being 1.5 the network has the same degree distribution as the BA model.
-Figure S.9. Dynamic evolution of networks created based on Principle 1.
-D.2. Principle 2: Triadic Closure
+Results for Principle 1 (preferential attachment): We display simulated networks comprising 200 nodes across different temperatures.
+For the degree-based simulations, node degree data $\{d_{j,t}:j\in V_{t}\}$ was provided ($V_{t}$ corresponds to the vertex set of the network $G_{t}$ at round $t$).
+With degree information only, the networks form more unrealistic star-like structures, diverging from scale-free configurations and more closely mirroring a core-periphery network structure.
+B.2. Principle 2: Triadic Closure
+Network Evolution
 We plot the evolution of the LLM-generated networks based on the triadic closure principle, together with the transitivity measure and the algebraic connectivity (which corresponds to the second-smallest eigenvalue of the graph Laplacian).
@@
 Figure S.10. Dynamic evolution of networks created based on Principle 2.
-Appendix E Feature Representations for Prompts
-Below, we give examples of the features used in the prompt presented in Algorithm 1. The features are formatted as a list of JSON objects which are provided to the prompt.
-E.1. Principle 1: Preferential Attachment
-For the structure-based information (Figure 1(a)) we have the following features:
-"name" : 0,
-"neighbors" : [5, 7, 1, 6]
-And for the degree-based information (Figure 1(b)) we have the following features:
-"name" : 0,
-"degree" : 4
-E.2. Principle 2: Triadic Closure
-For the structure-based information (Figure 2(a)) we have the following features:
-"name" : 0,
-"common_neighbors" : [5, 7, 1, 6]
-And for the degree-based information (Figure 2(b)) we have the following features:
-"name" : 0,
-"num_common_neighbors" : 4
-E.3. Principle 3: Homophily
-We have the following features:
-"name" : 0,
-"favorite_color" : "red",
-"hobby" : "hiking",
-"location" : "Boston"
-E.4. Principle 5: Small-World
-We have the following features:
-"name" : 0,
-"neighbors" : [5, 7, 1, 6]
-E.5. Real-World Data
-We have the following features:
-"name" : 0,
-"status" : "student",
-"major" : 10,
-"second major" : 93,
-"accommodation" : "house",
-"high_school" : 5,
-"graduation_year" : 2008
-We note that the initial Facebook100 dataset included gender information as a feature.
-We chose not to include the gender as one of the features as it has been shown that language models exhibit gender bias (Vig et al., 2020; Kotek et al., 2023; Bordia and Bowman, 2019).
+Simulations with the Number of Common Neighbors
+Instead of giving the neighborhood information, the simulations presented in Figure S.11 use the number of common neighbors. We observe behavior similar to Figure 2.
+Figure S.11.
+Results for Principle 2 (triadic closure).
+The figure shows the same networks as in Figure 2 with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative.
+Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test).
+The error bars correspond to 95% confidence intervals.
+Appendix C Chain-of-Thought Experiments
+# Task You are located in a school.
+Your task is to select a set of people to be friends with.
+# Profile Your profile is given below after chevrons: <PROFILE> { "name" : "Person 0", "favorite subject" : "Chemistry", "neighbors" : ["Person 3", "Person 432", "Person 4", "Person 3", "Person 32"] } </PROFILE> # Candidate Profiles The candidate profiles to be friends with are given below after chevrons: <PROFILES> [ { "name" : "Person 1", "favorite subject" : "Mathematics", "neighbors" : ["Person 3", "Person 4", "Person 23", "Person 65"] }, { "name" : "Person 33", "favorite subject" : "History", "neighbors" : ["Person 342", "Person 2", "Person 12"] }, ...
+] </PROFILES> # Output The output should be given a list of JSON objects with the following structure [ {{ "reason" : reason for selecting the person, "name" : name of the person you selected }}, ...
+] # Notes - The output must be a list of JSON objects ranked in the order of preference.
+- You can make at most 1 selection.
+ALGORITHM 3 Example prompt regarding social network data with Chain-of-Thought reasoning. Note that compared to Algorithm 1 the order of the fields name and reason in the output format is reversed.
+We experiment with Chain-of-Thought (CoT) reasoning (Wei et al., 2022).
+To induce CoT reasoning we ask the LLM agents to output the reason and then their choice (i.e.
+by reversing the order of reason and name in Algorithm 2).
+The resulting prompt can be found at Algorithm 3.
+In the following figures, we show the results from the same experiments as the ones we of the main text with the different that CoT is used.
+(a) Probability of connecting to top-$k$ nodes for different models, temperatures, and environments
+(b) Power law fits ($\hat{\gamma}$) and standard errors for different models, temperatures, and environments
+Figure S.12.
+Results for Principle 1 with CoT reasoning (preferential attachment) The multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$.
+Top: Probability of connecting to top-$k$-degree nodes for varying model (temperature is fixed to 1.0 and environment to baseline), temperature (model fixed to GPT-3.5 and environment to baseline) and environment (model fixed to GPT-3.5 and environment temperature to 1.5) for networks generated according to Principle 1 with $n=200$ nodes.
+Bottom: Power Law exponents and standard errors for varying model, temperature, and environment.
+(a) Probability of connecting to top-$k$ for different models, temperatures, and environments
+(b) Marginal transitivity ($D$) and probability of an edge within a community ($\hat{p}$) for different models, temperatures, and environments
+Figure S.13.
+Results for Principle 2 with CoT reasoning (triadic closure).
+Top: Probability of connecting to top-$k$ nodes (in terms of common neighbors) for varying model (temperature is fixed to 1.0 and environment to baseline), temperature (model fixed to GPT-4 Mini and environment to baseline) and environment (model fixed to GPT-4 Mini and environment temperature to 0.5) for networks generated according to Principle 2 ($n=50$, 10 simulations for each model, environment and temperature).
+Bottom: Marginal transitivity ($D$) and probability of an edge within a community ($\hat{p}$) for networks generated according to Principle 2 in different models, temperatures, and environments.
+(a) Assortativity and Louvain Modularity with different LLM models and environments
+Figure S.14.
+Results for Principle 3 (Homophily) and Principle 4 (Community structure due to homophily) with CoT reasoning.
+Top: Assortativities and Louvain modularity according to Principle 3 ($n=50$, 5 simulations for each row) in different environments (school, work, community) using different models.
+The statistical significance is $P<0.001$ for all t-tests (comparing with 0).
+(a) Regression plot for different models and environments for $\beta=0.25$ and $k=5$.
+Figure S.15.
+Fitted results for Principle 5 with CoT reasoning (small world).
+Regression plot for the relation $L\sim\log(n)$ for different LLM models for $\beta=0.25$ and $k=5$.
+The legend shows the effect size ($a$) and the $P$-value.
+(*: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.)